### v1

# Partial Search in a Frozen Network is Enough to Find a Strong Lottery Ticket

Abstract
Randomly initialized dense networks contain subnetworks that achieve high accuracy without weight learning—strong lottery tickets (SLTs). Recently, Gadhikar et al. (2023) demonstrated theoretically and experimentally that SLTs can also be found within a randomly pruned source network, thus reducing the SLT search space. However, this limits the search to SLTs that are even sparser than the source, leading to worse accuracy due to unintentionally high sparsity. This paper proposes a method that reduces the SLT search space by an arbitrary ratio that is independent of the desired SLT sparsity. A random subset of the initial weights is excluded from the search space by freezing it—i.e., by either permanently pruning them or locking them as a fixed part of the SLT. Indeed, the SLT existence in such a reduced search space is theoretically guaranteed by our subset-sum approximation with randomly frozen variables. In addition to reducing search space, the random freezing pattern can also be exploited to reduce model size in inference. Furthermore, experimental results show that the proposed method finds SLTs with better accuracy and model size trade-off than the SLTs obtained from dense or randomly pruned source networks. In particular, the SLT found in a frozen graph neural network achieves higher accuracy than its weight trained counterpart while reducing model size by 40.3×40.3$\times$40.3 ×.
Deep Learning, Strong Lottery Ticket, Model Compression
Figure 1: Freezing the source network by randomly pruning some parameters and locking others reduces the optimization targets for finding an SLT.
1 Introduction
The Strong Lottery Ticket Hypothesis (SLTH) conjectured the existence of subnetworks within a randomly weighted network—strong lottery tickets (SLTs)—that achieve comparable accuracy to trained dense networks (Zhou et al., 2019; Ramanujan et al., 2020; Malach et al., 2020). The existence of such subnetworks that do not require weight learning, illustrated in Figure 1(left), has been demonstrated experimentally (Zhou et al., 2019; Ramanujan et al., 2020; Huang et al., 2022; López García-Arias et al., 2023; Yeo et al., 2023) and proven theoretically for dense source networks (Malach et al., 2020; Orseau et al., 2020; Pensia et al., 2020; Diffenderfer & Kailkhura, 2021; da Cunha et al., 2021; Burkholz, 2022a; Burkholz et al., 2022; Ferbach et al., 2023; Gadhikar et al., 2023).
Since SLTs can be reconstructed from the random seed of the weights and a supermask (a simple binary pruning mask), potentially eliminating the need for storing and loading weights from memory, they offer a particularly advantageous opportunity for specialized inference hardware (Hirose et al., 2022). However, the existing algorithms for finding SLTs suffer from the same high computational cost as weight optimization, since they also employ backpropagation to update one parameter per neural connection (Ramanujan et al., 2020; Zhou et al., 2021; Diffenderfer & Kailkhura, 2021; Sreenivasan et al., 2022).
Recently, Gadhikar et al. (2023) demonstrated that this steep training cost can be alleviated by randomly pruning the source network at initialization, reducing the number of edges to be optimized (see Figure 1, center). However, this limits the SLT search space to the high sparsity region, which may not offer the optimal accuracy.
This paper introduces a novel method that reduces the search space by an arbitrary ratio independent of the desired SLT sparsity: in addition to random pruning at initialization, it also locks randomly chosen parameters at initialization to be a permanent part of the SLT (i.e., never pruned), as exemplified in Figure 1(right). Both the pruned and the locked parameters—the frozen parameters—are excluded from the search space, therefore reducing the training cost. Excluding them from the supermask also reduces the model size in inference, as the random freezing pattern can also be regenerated from its seed (Wimmer et al., 2020).
Moreover, our experiments show that SLTs found within frozen networks can achieve a better accuracy to model size trade-off than SLTs obtained from dense (Ramanujan et al., 2020; Sreenivasan et al., 2022) or sparse (Gadhikar et al., 2023) random networks. In addition to demonstrating this method experimentally, this paper also provides a mathematical analysis of the existence of SLTs within such partially frozen source networks.
In summary, this paper proposes a novel method that features the following contributions:
  • It reduces the SLT search space regardless of the desired SLT sparsity by random partial freezing (pruning and locking) of the source random network.
  • The proposed method is theoretically supported by our SLT existence result based on subset-sum approximation with randomly frozen variables.
  • Experimental results show that SLTs found in the frozen network achieve better accuracy-to-search space and accuracy-to-model size trade-offs than SLTs found within dense or sparse random networks.
Figure 2: Burkholz (2022a) proved that target network of depth $L$ can be stochastically approximated by an SLT of depth $L\!+\!1$ by approximating each target weights ($w_{T}$) by a subset-sum of the source weights. We extend this proof to frozen source networks by solving the subset-sum approximation with some source parameters that are either unavailable ($w^{\prime}_{2}$) or must be used for the subset-sum ($w^{\prime}_{0}$).
2 Preliminaries
After outlining the background on strong lottery tickets (SLTs) and their search algorithms, this section recapitulates the subset-sum approximation technique for the existence proof of SLTs (Pensia et al., 2020; da Cunha et al., 2021; Burkholz, 2022b, a) and its extension to sparse random networks (Gadhikar et al., 2023), which Section 3 expands to the existence of SLTs within frozen networks.
2.1 Strong Lottery Tickets in Dense Networks
SLTs (Zhou et al., 2019; Ramanujan et al., 2020; Malach et al., 2020) are subnetworks within a randomly weighted neural network that achieve high accuracy without any weight training. Compared with learned dense weight models, SLTs can be reconstructed from a small amount of information: since the random weights can be regenerated from their seed, it is only necessary to store the binary supermask and the seed (Hirose et al., 2022). Instead of updating weights, SLT search algorithms for deep neural networks (Ramanujan et al., 2020; Zhou et al., 2021; Sreenivasan et al., 2022) use backpropagation to update weight scores, which are then used to generate the supermask. For example, the Edge-Popup algorithm (Ramanujan et al., 2020; Sreenivasan et al., 2022), used in this paper, finds SLTs by applying a supermask generated by from the connections with the top-$k\%$ scores. The need to keep scores in addition to the weights, and the sorting of scores necessary for generating the supermask make the computational cost of Edge-Popup slightly higher than conventional weight training.
2.2 SLT Existence via Subset-Sum Approximation
Given a set of random variables and a target value, the subset-sum problem consists on finding a subset whose total value approximates the target. Lueker (1998) showed that such subset exists with high probability if the number of random variables is sufficiently large:
Lemma 2.1 (Subset-Sum Approximation (Lueker, 1998)).
Let $X_{1},...,X_{n}\sim U(-1,1)$ be independent, uniformly distributed random variables. Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by a subset-sum of $X_{i}$ if $n$ is sufficiently large.
Based on Lemma 2.1, previous works (Pensia et al., 2020; da Cunha et al., 2021; Burkholz, 2022b, a) showed that an SLT that approximates an arbitrary target network exists in a dense source network if it is wider and deeper than the target network. For example, Burkholz (2022a) proved that a source network with depth $L\!+\!1$ and larger width than the target network contains an SLT that can approximate the target network with depth $L$. As exemplified in Figure 2 (left), to approximate a weight $w_{T}$ of the target network, the SLT uses weights $w_{0},w_{1}$, and $w_{2}$ from the source network to construct the subset-sum. Therefore, the source network is required to be sufficiently wider than the target network, as stated in Lemma 2.1.
2.3 Strong Lottery Tickets in Sparse Networks
Recently, Gadhikar et al. (2023) revealed that SLTs also exist within sparse source networks, i.e., random networks that have been randomly pruned at initialization (see Figure 1, center). In addition to their theoretical contributions, their approach offered a practical advantage: since there are less parameters available from the beginning, the search space is reduced in proportion to the initial pruning ratio. This search space reduction can be exploited for reducing the computational cost of training.
2.4 SLT Existence in Sparse Networks
To prove the existence of SLTs within sparse networks, Gadhikar et al. (2023) extended the subset-sum approximation (Lemma 2.1) to the situation where randomly chosen variables are permanently pruned at initialization:
Lemma 2.2 (Subset-Sum Approximation in Sparse Networks (Gadhikar et al., 2023)).
Let $X_{1},...,X_{n}$ be as in Lemma 2.1, and $M_{1},...,M_{n}\sim Ber(p)$ be independent, Bernoulli distributed random variables with $p\in(0,1)$.Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by a subset-sum of $M_{i}X_{i}$ if $n$ is sufficiently large.
By applying this extended lemma instead of Lemma 2.1 to the SLT existence theorem presented by Burkholz (2022a), Gadhikar et al. (2023) proved the SLT existence in sparse networks as the following theorem:
Theorem 2.3 (SLT Existence in Sparse Networks (Gadhikar et al., 2023)).
Let a target network $f_{T}$ with depth $L$ and a sparse source network $f_{S}$ with depth $L\!+\!1$. Assume that the source network is randomly pruned with pruning ratio $p_{l}$ for each $l$-th layer. Also assume that these networks have ReLU activation function and initialized by using a uniform distribution $U[-1,1]$. Then, except with exponentially small probability, a subnetwork $f_{SLT}$ exists in the sparse source network $f_{S}$ such that $f_{SLT}$ approximates the target network $f_{T}$ if the width of $f_{S}$ is sufficiently large.
Thus, by extending the subset-sum approximation lemma, the conventional SLT existence proof can be easily extended to various network settings.
3 Strong Lottery Tickets in Frozen Networks
Although pruning at initialization provides a search space reduction, it also imposes a limitation on the sparsity range, as it can only search for SLTs with a sparsity higher than the initial pruning ratio. For example, although an initial pruning ratio of $90\%$ vastly reduces the search space, it limits the search to only SLTs sparser than $90\%$. Therefore, this reduction in search space is incompatible with exploring the whole sparsity range for optimal accuracy.
This paper proposes a novel method that reduces the search space regardless of the desired SLT sparsity to capitalize on the training cost reduction without sacrificing accuracy. We explore the existence and performance of SLTs within a frozen source network, i.e., a random network that, in addition to randomly pruned parameters, has randomly chosen parameters forced to be permanent (i.e., never pruned) part of the SLT—locked parameters (see Figure 1, right). Since freezing is performed at initialization, before the SLT search, frozen parameters do not have a score and are not eligible for exploration, thus reducing the search space. Locking allows to further reduce the parameter search without limiting the initial pruning ratio, allowing to reduce the search space even in the case of dense source networks.
This section first provides a theoretical analysis on the existence of SLTs in such a frozen network by developing a new extension of the subset-sum approximation introduced in Section 2. Then, based on this analysis, it proposes partial SLT search, a novel method for finding such SLTs with a reduced search space. In preparation for the evaluation experiments in Section 4, we also perform a preliminary investigation of the optimal settings of the proposed method.
3.1 SLT Existence in Frozen Networks
Here we provide a theoretical result that indicates that an SLT capable of approximating a target network exist in a frozen network. This is proved by extending the subset-sum approximation lemma (Lemma 2.1) to the case where some parameters are locked (for detailed proof see Appendix A).
Lemma 3.1 (Subset-Sum Approximation in Randomly Locked Networks (Lemma A.2)).
Let $X_{1},...,X_{n}$ be as in Lemma 2.1, and $M^{\prime}_{1},...,M^{\prime}_{n}\sim Ber(q)$ be independent, Bernoulli distributed random variables with $q\in(0,1)$.Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by the sum of $\sum_{i=1}^{n}M^{\prime}_{i}X_{i}$ and a subset-sum of $(1-M^{\prime}_{i})X_{i}$ if $n$ is sufficiently large.
Then, combining Lemma 3.1 with Lemma 2.2 further extends the subset-sum approximation to the situation where some random variables are frozen (i.e., pruned or locked).
Lemma 3.2 (Subset-Sum Approximation in Frozen Networks (Lemma A.3)).
Let $X_{1},...,X_{n}$ be as in Lemma 2.1, $M_{1},...,M_{n}\sim Ber(p)$ be as in Lemma 2.2, and $M^{\prime}_{1},...,M^{\prime}_{n}\sim Ber(q)$ be as in Lemma 3.1. Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by the sum of $\sum_{i=1}^{n}M_{i}M^{\prime}_{i}X_{i}$ and a subset-sum of $M_{i}(1-M^{\prime}_{i})X_{i}$ if $n$ is sufficiently large.
Finally, by applying Lemma 3.2 to the Theorem 2.5 in Gadhikar et al. (2023) instead of Lemma 2.2, it follows that an SLT, which approximates a target network, exists in frozen networks.
Theorem 3.3 (SLT Existence in Frozen Networks (Theorem A.4)).
Let a target network $f_{T}$ with depth $L$ and a partially frozen source network $f_{S}$ with depth $L\!+\!1$ be given. Assume that the source network is randomly frozen with pruning ratio $p_{l}$ and locking ratio $q_{l}$ for each $l$-th layer. Also assume that these networks have ReLU activation function and initialized by using a uniform distribution $U[-1,1]$. Then, except with exponentially small probability, a subnetwork $f_{SLT}$ exists in the frozen source network $f_{S}$ such that $f_{SLT}$ approximates the target network $f_{T}$ if the width of $f_{S}$ is sufficiently large.
Figure 3: Pruning and locking set the bounds of the search space, whose optimal position is investigated.
Figure 3: Pruning and locking set the bounds of the search space, whose optimal position is investigated.
Figure 4: Comparison of different prune:lock proportions of a $80\%$ freezing ratio using a Conv6 on CIFAR-1010{10}10.
Figure 5: Impact of the freezing ratio on different architectures. Pruning and locking ratios are set following Section 3.3.
3.2 Partial SLT Search in Frozen Networks
This paper proposes a novel method for searching SLTs within frozen networks: partial SLT search. It only searches in part of the source network, as it has been pruned at initialization; and it only searches for part of the final SLT, as some of its parameters are predefined by locking.
Figure 4 illustrates how network freezing reduces the search space. The total amount of frozen parameters is determined by a global freezing ratio $F_{r}$, which is the sum of the respective pruning ratio $P_{r}$ and locking ratio $L_{r}$. Therefore, the SLT search space $S$ is formed by the $1\!-\!F_{r}$ non-frozen parameters of the original dense network. As explained previously, it is not possible to search for an SLT with a lower sparsity (denser) than the source network. Consequently, the pruning ratio sets the lower bound of the search space. On the other hand, it is not possible to search for an SLT that prunes more parameters than available, so the locking ratio sets the upper bound of the search space. (Specifically, the upper bound is given by $1\!-\!F_{r}$) Thus, these ratios allow to freely control the size and position of the search space over the SLT sparsity spectrum.
Frozen network construction:
We assume that the $l$-th layer of network with depth $L$ has weights $\bm{w}^{(l)}\in\mathbb{R}^{n^{(l)}}$. The number of frozen parameters is determined by the layer-wise freezing ratios $p_{f}^{(1)},...,p_{f}^{(L)}$, where each $p_{f}^{(l)}$ is defined as the sum of the layer-wise pruning ratio $p_{p}^{(l)}$ and the layer-wise locking ratio $p_{l}^{(l)}$, and $p_{f}^{(l)},p_{p}^{(l)},p_{l}^{(l)}\in[0,1]$.
To prune $p_{f}^{(l)}n^{(l)}$ parameters and lock $p_{l}^{(l)}n^{(l)}$ parameters, we generate two random masks, a pruning mask $\bm{m}^{(l)}_{p}\in\{0,1\}^{n^{(l)}}$ and a locking mask $\bm{m}^{(l)}_{l}\in\{0,1\}^{n^{(l)}}$, so that they satisfy $||1-\bm{m}^{(l)}_{p}||=p_{p}^{(l)}n^{(l)}$, $||\bm{m}^{(l)}_{l}||=p_{l}^{(l)}n^{(l)}$, and $(1-\bm{m}^{(l)}_{p})\cdot\bm{m}^{(l)}_{l}=0$ (i.e., we require pruning and locking to be implemented without overlap). The layer-wise weights frozen with these masks are calculated as
$\bm{w}^{(l)}_{f}$ $:=(\bm{m}_{p}^{(l)}\odot(1-\bm{m}_{l}^{(l)})+\bm{m}_{l}^{(l)})\odot\bm{w}^{(l)}.$ (1)
These masks are fixed during training, and we search for SLTs only in the parts where $\bm{m}_{p}^{(l)}\odot(1-\bm{m}_{l}^{(l)})$ is one.
Setting the layer-wise ratios:
Our method considers the following two existing strategies for determining the layer-wise pruning ratio from the desired global pruning ratio of the source network:
  • Edge Per Layer (EPL): Each layer’s pruning ratio is set so that they all have the same number of remaining weights (Price & Tanner, 2021; Gadhikar et al., 2023).
  • Erdős-Rényi-Kernel (ERK): The pruning ratio of layer $l$ is proportional to the scale $(C_{in}^{(l)}+C_{out}^{(l)}+k_{h}^{(l)}+k_{w}^{(l)})/(C_{in}^{(l)}\times C_{out}^{(l)}\times k_{h}^{(l)}\times k_{w}^{(l)})$, where $C_{in}^{(l)}$, $C_{out}^{(l)}$, $k_{h}^{(l)}$, and $k_{w}^{(l)}$ denote input channels, output channels, kernel height, and kernel width of the layer $l$, respectively (Evci et al., 2020).
The same strategy is employed to determine the layer-wise pruning and freezing ratios from their respective global ratios, and then locking ratios are calculated as the difference between the corresponding freezing and pruning ratios.
Freezing mask compression:
The freezing pattern can be encoded during inference as a ternary mask—a freezing mask—that indicates whether a parameter is pruned, locked, or part of the supermask. For example, encoding pruning as $-1$, locking as $+1$, and supermask inclusion as 00, the layer-wise freezing mask can be encoded as $\bm{m}^{(l)}_{l}+(\bm{m}^{(l)}_{p}-1)$. Since this freezing mask is also random and fixed, it can be regenerated from its seed and ratios, similarly to the random weights. Furthermore, the supermask size is reduced by excluding from it the frozen parameters, so the SLTs found by this method can be compressed in inference to an even smaller memory size than those produced by the existing SLT literature (Hirose et al., 2022; Okoshi et al., 2022; López García-Arias et al., 2023).
3.3 Optimal Positioning of the Search Space
Here we perform a preliminary investigation of the optimal positioning for each given desired SLT sparsity by exploring different pruning:locking proportions with a fixed freezing ratio of the network. Figure 4 explores different configurations of a $80\%$ freezing ratio—a search space size of $20\%$ of the original dense source network—on a Conv6 network and compares the performance of the found SLTs on CIFAR-1010{10}10. As expected, the best-performing SLT of each prune:lock configuration is found at the center of the search space, which maximizes the number of candidate subnetworks. In other words, within a search space size $S$, the optimal SLT sparsity is $k=P_{r}+S/2$, where $P_{r}=\frac{\sum_{l}p^{(l)}_{p}n^{(l)}}{\sum_{k}n^{(k)}}$ is the global pruning ratio of the network. Conversely, for a given freezing ratio $F_{r}$ of the network and a desired SLT sparsity $k$, the optimal search space position is set by the pruning ratio $P_{r}=k-(1-F_{r})/2$ and the corresponding locking ratio $L_{r}=F_{r}-P_{r}$. In the cases where this would result in $P_{r}\!<\!0$ or $L_{r}\!<\!0$, we choose a best-effort approach that keeps the desired freezing ratio and sets the search space bounds to $P_{r}\!=\!0$ or $L_{r}\!=\!0$, respectively.
Additionally, Figure 4 compares the two strategies for setting ratios considered in Section 3.2, showing that EPL outperforms ERK in all cases. Consequently, hereafter the proposed method sets the global ratios in order to position the search space center as close as possible to the desired SLT sparsity, and then sets the layer-wise ratios using EPL.
Figure 6: Freezing achieves better accuracy to search space trade-off compared to sparse ($\bullet$) or dense ($\bullet$) source networks. SLT sparsity set according to highest accuracy in Figure 5 with dense source networks.
4 Experiments
This section demonstrates that partial SLT search reduces the search space in a broad range of situations by evaluating it on image classification and node classification. We evaluate the impact of the freezing ratio on various network architectures, identifying three scenarios. Then, we explore trade-offs between accuracy, search space size, and model memory size for different network widths.
4.1 Experimental Settings
The proposed partial SLT search is evaluated on image classification using the CIFAR-1010{10}10 (Krizhevsky, 2009) dataset, and on node classification using the OGBN-Arxiv (Hu et al., 2020) dataset. CIFAR-1010{10}10 train data is split into training and validation sets with a 4:1 ratio, while for OGBN-Arxiv we use the default set split. We test the models with the best validation accuracy, and report the mean of three experiment repetitions.
For image classification we employ the VGG-like Conv6 (Simonyan & Zisserman, 2014; Ramanujan et al., 2020) and ResNet-1818{18}18 (He et al., 2016) architectures, and for graph node classification the GIN (Xu et al., 2019) architecture constructed with 4 layers modified by Huang et al. (2022), all implemented with no learned biases. ResNet-1818{18}18 and GIN use non-affine Batch Normalization (Ioffe & Szegedy, 2015), while Conv6 has no normalization layers. Random weights are initialized with the Kaiming Uniform distribution, while weight scores are initialized with the Kaiming Normal distribution (He et al., 2015).
SLTs are searched using an extension of Edge-Popup (Ramanujan et al., 2020) that enforces the desired sparsity globally instead of per-layer (Sreenivasan et al., 2022). On CIFAR-1010{10}10, scores are optimized for $100$ epochs using stochastic gradient descent (SGD) with momentum $0.9$, batch size $128$, weight decay $0.0001$, and initial learning rates of $0.01$ and $0.1$ for Conv6 and ResNet-1818{18}18, respectively. On OBGN-Arxiv, scores are optimized for $400$ epochs using AdamW (Loshchilov & Hutter, 2019) with weight decay $0.0001$ and initial learning rate $0.01$. All experiments use cosine learning rate decay (Loshchilov & Hutter, 2017).
Figure 7: Freezing achieves better accuracy to memory size trade-off compared to sparse ($\bullet$) or dense ($\bullet$) source networks.
4.2 Reducing the Search Space with Freezing
An experiment is first conducted to investigate the effect of the freezing ratio of the network on SLT accuracy. We identify three scenarios, represented in Figure 5.
In the cases where the optimal SLT sparsity is found at intermediate sparsity—e.g., around $50\%$ for Conv6 in LABEL:fig:sparsity_acc_conv6—pruning and locking can be applied with equally high ratios. Results show that the search space can be reduced by $40\%$ with small impact on accuracy, and by $70\%$ with still a moderate accuracy drop of $5$ points.
Applying the much larger ResNet-1818{18}18 to the same task results in much stronger overparametrization. Therefore, optimal SLTs are found in the higher sparsity range, as revealed by LABEL:fig:sparsity_acc_resnet18, benefiting of much higher pruning than locking. Even in this scenario, $90\%$ sparse SLTs found in frozen networks achieve similar accuracy ($<1$ point difference) while reducing the search space by as much as $80\%$.
As an example of the scenario where optimal SLT sparsities are found in the denser range, benefitting from a higher locking ratio, we evaluate GIN in LABEL:fig:sparsity_acc_gin. Compared with the best-performing SLT found in the dense GIN, of $70.1\%$ accuracy and $20\%$ sparsity, by freezing $50\%$ of the search space our method finds better performing SLTs of $70.2\%$ accuracy with $40\%$ sparsity.
These results suggest that parameter freezing at an appropriate ratio has the effect of eliminating local optimal solutions from the search space. It is known that Edge-Popup struggles to find the SLTs of the global optimal solution (Fischer & Burkholz, 2022). While searching for SLTs in a dense network leads to convergence to a local optimal solution, a moderate reduction of the search space may reduce the number of less accurate local optimal solutions and facilitate convergence to a more accurate local optimal solution within the reduced search space. In other words, we conjecture that if the network is properly frozen, the local optimal solution for the reduced search space is close to the global optimal solution of the entire search space.
4.3 Accuracy to Search Space Trade-Off
Section 3 sets as a requisite of the subset-sum proof that the source network has a sufficiently larger width than the target network. Figure 5 explores varying the width of the source network to analyze its impact on accuracy and search space size. SLT sparsity is fixed to that of the best performing SLT found in a dense source network in Figure 5: $50\%$ in Conv6, $90\%$ in ResNet-1818{18}18, and $20\%$ in GIN.
Compared to the dense source network, the results show in all cases that, when the frozen network is twice as wide, the found SLTs for a given sparsity achieve similar or higher performance with a similar or smaller search space size. That is, the propsed method improves the accuracy to search space size trade-off. In the case of GIN in LABEL:fig:search_space_acc_gnn, the SLT found in a 2×2$\times$2 × wider frozen source even achieves higher accuracy than those found in any dense source network.
As shown in LABEL:fig:search_space_acc_conv6 and LABEL:fig:search_space_acc_gnn, in the scenarios that only allow for medium to low pruning, there is a limit to the number of parameters that can be pruned for a given SLT sparsity, and only aggressive pruning leads to less accurate SLTs. By additionally locking parameters, our method reduces the search space further and finds SLTs of higher accuracy than those within sparse or dense networks for a similar search space size. Pruning only improves the trade-off in the scenario that allows higher pruning than locking, as shown in LABEL:fig:search_space_acc_resnet18, while our method improves the accuracy to search space trade-off in all scenarios.
4.4 Accuracy to Model Memory Size Trade-Off
The freezing mask compression scheme proposed in Section 3.2 allows to reduce the memory model size during inference by regenerating both the random weights and the freezing mask with a random number generator (RNG). Here we consider this compression and investigate the accuracy to model size trade-off offered by the SLTs found by the proposed method. Furthermore, considering the practical application of partial SLT search on specialized hardware, we also compare the Kaiming Uniform (KU) weight initialization used so far with the binary weights provided by the Signed Kaiming Constant (SKC) initialization (Ramanujan et al., 2020), which can be exploited for reduced computational cost in neural engines (Hirose et al., 2022). For SKC we scale weights by $1/\sqrt{1-k_{l}}$, where $k_{l}$ is the sparsity of each layer, as proposed by Ramanujan et al. (2020).
As illustrated in Figure 7, compared to the SLTs found in dense or sparse source networks, SLTs found by our method achieve similar or higher accuracy for similar or smaller model size, thus improving the accuracy to model size trade-off in all scenarios. In other words, partial SLT search not only reduces the search space during training, but also reduces the model size during inference, further improving performance over conventional SLT search methods.
Empirically, it is known that SLTs within a SKC initialized dense network achieve better performance than with continuous random weights (Ramanujan et al., 2020; Okoshi et al., 2022; López García-Arias et al., 2023; Yan et al., 2023). Our results show that such a trend can also be observed in frozen source networks. Nonetheless, in all source networks, we find that this improvement is smaller the wider the source network is, suggesting that the requirement for source networks of larger width is weaker in the case of binary weights.
Table 1: Comparison between trained-weight networks and SLTs found in dense, sparse, and frozen networks.
Conv6 & CIFAR-1010{10}10
Method
(Source Network)
Weight
Initialization
Sparsity
Pruning
Ratio [%]
Locking
Ratio [%]
Top-1 Test
Acc. [%]
Search Space
[# Parameters]
Model Size
[MB]
Weight Training KU - - - $87.1$ $2.26$ M $8.63$
SLT (Dense) SKC $50$ 00 00 $86.2$ $2.26$ M $0.27$
SLT (Sparse) SKC $50$ $45$ 00 $66.7$ $1.24$ M $0.15$
SLT (Frozen) SKC $50$ $25$ $25$ $84.8$ $1.13$ M $0.13$
ResNet-${18}$ CIFAR-1010{10}10
Weight Training KU - - - $92.5$ $11.16$ M $42.63$
SLT (Dense) SKC $90$ 00 00 $91.0$ $11.16$ M $1.37$
SLT (Sparse) SKC $90$ $85$ 00 $89.9$ $1.67$ M $0.24$
SLT (Frozen) SKC $90$ $82.5$ $2.5$ $89.8$ $1.67$ M $0.24$
SLT (Frozen) SKC $90$ $85$ $5$ $88.9$ $1.12$ M $0.17$
GIN & OGBN-Arxiv
Weight Training KU - - - $70.1$ $0.38$ M $1.452$
SLT (Dense) SKC $20$ 00 00 $70.1$ $0.38$ M $0.054$
SLT (Sparse) SKC $20$ $15$ 00 $63.1$ $0.32$ M $0.047$
SLT (Frozen) SKC $20$ 00 $40$ $70.6$ $0.23$ M $0.036$
4.5 Result Analysis
Table 1 summarizes the presented results and compares the proposed method to weight training and SLT search in dense and sparse source networks for the same network architecture and SLT sparsity.
In the case of Conv6, compared to weight training our method provides reductions of 2×2$\times$2 × and 66.4×66.4$\times$66.4 × of the search space and model size, respectively, in exchange of a small accuracy drop. Compared to the SLT found in the sparse source network, the SLT found by partial search achieves much higher accuracy.
[…]
Interestingly, the SLT found in the frozen GIN model achieves higher accuracy than the trained-weight network, even though the model size and search space are reduced by 40.3×40.3$\times$40.3 × and 1.7×1.7$\times$1.7 ×, respectively. The SLT found within a dense network also achieves comparable accuracy to the trained-weight network, but the SLT found by our method performs better and is smaller. Generally, graph neural networks (GNNs) suffer from a generalization performance degradation due to the over-smoothing problem (Li et al., 2018). Searching for SLTs within dense GNNs has been shown to mitigate the over-smoothing problem (Huang et al., 2022) and achieve higher accuracy. As our method is even more accurate than dense networks, it is possible that parameter freezing offers an even stronger solution to the over-smoothing problem.
5 Conclusion
This paper capitalizes on the fact that pruning a randomly weighted network at initialization reduces the SLT search space, but identifies that it limits the search to a suboptimal sparsity region. This problem is tackled by freezing (i.e., pruning or locking) some parameters at initialization, excluding them from the search. Freezing allows to search for SLTs efficiently in the optimal sparsity range, while further reducing search space during training and model size during inference. Experimental results show that SLTs found in frozen networks improve the accuracy to search space and accuracy to model size trade-offs compared to SLTs found in dense (Ramanujan et al., 2020; Sreenivasan et al., 2022) or sparse source networks (Gadhikar et al., 2023). Interestingly, although for weight training random parameter locking has only been found useful for reducing accuracy degradation in network compression (Zhou et al., 2019; Wimmer et al., 2020), we identify a scenario in SLT training where it can be used for raising accuracy, even outperforming weight-trained networks. Our method can be interpreted as being capable of generating more useful SLT information from a random seed than previous methods, and as a result, it offers an opportunity for reducing off-chip memory access in specialized SLT accelerators (Hirose et al., 2022). We hope that this work will serve as a base for raising further the efficiency of SLT training and inference.
Impact Statements
This paper presents work whose goal is to advance the field of machine learning. Among the many potential societal consequences of our work, we highlight its potential to reduce the computational cost of inference and training of neural networks, and thus their energy consumption footprint.
Appendix A Proof for Strong Lottery Tickets (SLTs) Existence in Frozen Networks
This section describes the proofs of the lemma and theorem introduced in the manuscript. For reference, we present a theorem presented by Gadhikar et al. (2023) in advance.
Theorem A.1 (SLT Existence in Sparse Networks (Gadhikar et al., 2023)).
Let input data $\mathcal{D}$ , a target network $f_{T}$ with depth $L$, and a source network $f_{S}$ with depth $L+1$, parameters $\bm{\theta}_{S}$, and edge probabilities $p_{l}$ be given. Assume that these networks have ReLU activation function, and each element of $\bm{\theta}_{S}$ is uniformly distributed over $[-1,1]$. Then, with probability at least $1-\delta$, there exists a binary mask $\bm{m}_{S}$ so that each output component $i$ is approximated as $\max_{\bm{x}\in\mathcal{D}}\left\|f_{T,i}(\bm{x})-f_{S,i}(\bm{x};\bm{\theta}_{S}\odot\bm{m}_{S})\right\|<\epsilon$ if
$n_{S,l}\geq C\frac{n_{T,l}}{\log(1/1-p_{l+1})}\log\left(\frac{1}{\min\{\epsilon_{l},\delta/\rho\}}\right)$ (2)
for $l\geq 1$,where
$\epsilon_{l}$ $\coloneqq\frac{\epsilon}{n_{T,L}L}\left[(1+B_{l-1}\left(1+\frac{\epsilon}{L}\right)\prod_{k=l+1}^{L-1}\left(\left\|\bm{\theta}_{T}^{(l)}\right\|_{\infty}+\frac{\epsilon}{L}\right)\right]^{-1},$ (3)
$B_{l}$ $\coloneqq\sup_{\bm{x}\in\mathcal{D}}\left\|\bm{x}_{T}^{(l)}\right\|_{1},$ (4)
$\rho$ $=\frac{CN_{T}^{1+\gamma}}{\log\left(1/(1-\min_{l}\{p_{l}\})\right)^{1+\gamma}}\log\left(\frac{1}{\min\{\min_{l}\{\epsilon_{l}\},\delta\}}\right)$ (5)
for any $\gamma\geq 0$. Moreover, we need to satisfy
$n_{S,0}\geq Cd\log\left(\frac{1}{\min\left\{\epsilon_{1},\delta/\rho\right\}}\right),$ (6)
where $C>0$ denotes a generic constant that is independent of $n_{T,l},L,p_{l},\delta$, and $\epsilon$.
We first extend the lemma of the subset-sum approximation shown by Lueker (1998) to the case where some random variables are always included in the subset-sum.
Lemma A.2 (Subset-Sum Approximation in Randomly Locked Networks).
Let $X_{1},\cdots,X_{n}\sim U(-1,1)$ be independent, uniformly distributed random variables, and $M_{1},\cdots,M_{n}\sim\mathrm{Ber}(q)$ be independent, Bernoulli distributed random variables with $q\in(0,1)$. Let $\varepsilon,\delta>0$. Then, with probability at least $1-\delta$, for any $z\in[-1,1]$, there exists indices $I\subset\{1,\cdots,n\}$ such that $|z-\sum_{i=1}^{n}M_{i}X_{i}-\sum_{i\in I}(1-M_{i})X_{i}|\leq\varepsilon$ if
$n\geq C\log\left(\frac{8}{\varepsilon\delta}\right),$ (7)
where $C>0$ is a constant.
Proof.
Let $M_{1},\cdots,M_{n}\sim\mathrm{Ber}(q)$ and $m:=\sum_{i}M_{i}$. By Hoeffding’s inequality, we have
$\mathbb{P}\left(|m-qn|\leq\varepsilon_{M}n\right)\geq 1-2\exp\left(-\frac{\varepsilon_{M}^{2}n}{2q(1-q)}\right)$ (8)
for $\varepsilon_{M}>0$. Thus, if we set $n\geq\frac{2q(1-q)\log(2/\delta)}{\varepsilon_{M}^{2}}$ and $\varepsilon_{M}:=\frac{\min(q,1-q)}{2}$, we have
$(2q-\beta)n\leq m\leq\beta n$ (9)
with $\beta:=q+\varepsilon_{M}=\min(\frac{3q}{2},\frac{1+q}{2})\in(0,1)$, with probability at least $1-\delta$. In particular, the number of non-vanishing terms in the sum $\sum_{i\in I}(1-M_{i})X_{i}$ is $n-m\geq(1-\beta)n$ as long as each $X_{i}$ is non-zero.
Now fix $M_{1},\cdots,M_{n}\sim\mathrm{Ber}(q)$ with $(2q-\beta)n\leq m\leq\beta n$. The goal is to approximate $z\in[-1,1]$ and $\sum_{i=1}^{n}M_{i}X_{i}$ by subset sum from $\{(1-M_{i})X_{i}:M_{i}=0\}$. For simplicity, we split it into two parts:
$\{(1-M_{i})X_{i}:M_{i}=0\}=\{X^{\prime}_{1},\cdots,X^{\prime}_{n_{1}}\}\cup\{X^{\prime\prime}_{1},\cdots,X^{\prime\prime}_{n_{2}}\},$ (10)
where the former part is used for approximating $z\in[-1,1]$ and the latter part for approximating $\sum_{i=1}^{n}M_{i}X_{i}$.
To approximate $z\in[-1,1]$, we can directly apply Corollary 2.5 from Lueker (1998):
$\mathbb{P}\left(\forall z\in\left[-1,1\right],\exists I\subset\{1,\cdots,n_{1}\}\text{ s.t. }|z-\sum_{i\in I}X^{\prime}_{i}|\geq\varepsilon\right)\geq 1-\delta$ (11)
whenever $n_{1}\geq C\log(2/\varepsilon\delta)$.
To approximate $\sum_{i=1}^{n}M_{i}X_{i}$, we have to evaluate its norm. By Hoeffding’s inequality on $X_{i}$’s with $M_{i}\not=0$, we have
$\mathbb{P}\Big{(}\Big{|}\sum_{i=1}^{n}M_{i}X_{i}\Big{|}\leq\alpha m\Big{)}\geq 1-2\exp\left(-\frac{3\alpha^{2}m}{2}\right),$ (12)
for any fixed $\alpha>0$, whose value will be specified later. Thus $\Big{|}\sum_{i=1}^{n}M_{i}X_{i}\Big{|}\leq\alpha m\leq\alpha\beta n$ holds with probability at least $1-\delta$ whenever $m\geq\frac{2\log(2/\delta)}{3\alpha^{2}}$. Since $m\leq\beta n$ holds, $n\geq\frac{2\log(2/\delta)}{3\alpha^{2}\beta}$ is enough.
From the proof of Corollary 3.1 in Lueker (1998), for any $\gamma\in(0,\frac{1}{4})$, we know that
$\mathbb{P}\left(\forall z\in\left[-\gamma n_{2},\gamma n_{2}\right],\exists I\subset\{1,\cdots,n_{2}\}\text{ s.t. }|z-\sum_{i\in I}X^{\prime\prime}_{i}|\geq\varepsilon\right)$ (13)
$\geq 1-\delta-2\exp\left(-\frac{(1-4\gamma)^{2}n_{2}}{64}\right)$ (14)
whenever $n_{2}\geq C\log(\frac{2}{\varepsilon\delta})$. Thus if
$n_{2}\geq\max\left(C\log(\frac{2}{\varepsilon\delta}),\frac{C^{\prime}}{(1-4\gamma)^{2}}\log(\frac{2}{\delta})\right)$ (15)
holds, we can approximate any $z\in[-\gamma n_{2},\gamma n_{2}]$ by subset sum of $X^{\prime\prime}_{1},\cdots,X^{\prime\prime}_{n_{2}}$ with probability $1-2\delta$.
Now we assume $(\alpha\beta/\gamma)n\leq n_{2}$. For example, if we set $\alpha=\frac{1}{10}$ and $\gamma=\frac{1}{5}$, the assumption is satisfied when we split $\{(1-M_{i})X_{i}\}$ by $n_{1}=n_{2}=\frac{n}{2}$ in (10). Then the approximation (13) can be applied to $z=\sum_{i=1}^{n}M_{i}X_{i}$ with high probability since $|\sum_{i=1}^{n}M_{i}X_{i}|\leq\alpha m\leq\alpha\beta n\leq\gamma n_{2}$ holds.
By combining (11), (12), (15) with the fact that $n=m+n_{1}+n_{2}$, and replacing $\delta$ with $\delta/4$, we obtain the desired results. ∎
By similar arguments as in Lueker (1998) or Pensia et al. (2020), we can easily generalize Lemma A.2 to the case where the distribution followed by $X_{i}$ contains the uniform distribution. Also, by combining our proof and the proof of Lemma 2.4 in Gadhikar et al. (2023), we can prove the following extension of Lemma A.2:
Lemma A.3 (Subset-Sum Approximation in Frozen Networks.).
Let $X_{1},...,X_{n}$ be independent, uniformly distributed random variables so that $X\sim U(-1,1)$, $M_{1},...,M_{n}$ be independent, Bernoulli distributed random variables so that $M_{i}\sim Ber(p)$ for $p\in(0,1)$, and $M^{\prime}_{1},...,M^{\prime}_{n}$ be independent, Bernoulli distributed random variables so that $M^{\prime}_{i}\sim Ber(q)$ for $q\in(0,1)$. Let $\epsilon,\delta\in(0,1)$ given. Then for any $z\in[-1,1]$ there exists a subset $I\subseteq\{1,...,n\}$ so that with probability at least $1-\delta$ we have $\left|z-\sum_{i=1}^{n}M_{i}M^{\prime}_{i}X_{i}-\sum_{i\in I}M_{i}\left(1-M^{\prime}_{i}\right)X_{i}\right|\leq\epsilon$ if
$n\geq C\log\left(\frac{C^{\prime}}{\varepsilon\delta}\right),$ (16)
where $C,C^{\prime}>0$ is a constant.
Theorem A.4 (SLT Existence in Frozen Networks).
Let $\mathcal{D}$ be input data, $f_{T}$ be a target network with depth $L$, and $f_{S}$ be a source network with depth $L+1$, $\bm{\theta}_{S}$ be a parameter of $f_{S}$, and $p_{l},q_{l}\in(0,1)$. Assume that these networks have ReLU activation function, and each element of $\bm{\theta}_{S}$ is uniformly distributed over $[-1,1]$. Also assume that $\bm{\theta}_{S}$ is randomly pruned and locked with pruning ratio $p_{l}$ and locking ratio $q_{l}$ for each $l$-th layer. Then, with probability at least $1-\delta$, there exists a binary mask $\bm{m}_{S}$ so that each output component $i$ is approximated as $\max_{\bm{x}\in\mathcal{D}}\left\|f_{T,i}(\bm{x})-f_{S,i}(\bm{x};\bm{\theta}_{S}\odot\bm{m}_{S})\right\|<\epsilon$ if
$n_{S,l}$ $\geq C_{1}\log\left(\frac{C_{2}}{\min\{\epsilon_{l},\delta/\rho\}}\right)n_{T,l},$ $(l\geq 1)$ (17)
$n_{S,0}$ $\geq C_{3}\log\left(\frac{C_{4}}{\min\left\{\epsilon_{1},\delta/\rho\right\}}\right)d$ $(l=0)$ (18)
where $C_{i}>0$ are constants that include $p_{l},q_{l}$, and $\rho,\epsilon_{l}$ are as defined in Theorem A.1.
Proof.
By replacing Lemma 2.4 in Gadhikar et al. (2023) with our Lemma A.3, we obtain the desired result as an extension of Theorem 2.5 in Gadhikar et al. (2023). ∎

### v2

@@
 Recently, Gadhikar et al.
-(2023) demonstrated theoretically and experimentally that SLTs can also be found within a randomly pruned source network, thus reducing the SLT search space.
+[10] demonstrated that SLTs can also be found within a randomly pruned source network, thus reducing the SLT search space.
 However, this limits the search to SLTs that are even sparser than the source, leading to worse accuracy due to unintentionally high sparsity.
-This paper proposes a method that reduces the SLT search space by an arbitrary ratio that is independent of the desired SLT sparsity.
+This paper proposes a method that reduces the SLT search space by an arbitrary ratio independent of the desired SLT sparsity.
 A random subset of the initial weights is excluded from the search space by freezing it—i.e., by either permanently pruning them or locking them as a fixed part of the SLT.
-Indeed, the SLT existence in such a reduced search space is theoretically guaranteed by our subset-sum approximation with randomly frozen variables.
-In addition to reducing search space, the random freezing pattern can also be exploited to reduce model size in inference.
-Furthermore, experimental results show that the proposed method finds SLTs with better accuracy and model size trade-off than the SLTs obtained from dense or randomly pruned source networks.
-In particular, the SLT found in a frozen graph neural network achieves higher accuracy than its weight trained counterpart while reducing model size by 40.3×40.3$\times$40.3 ×.
-Deep Learning, Strong Lottery Ticket, Model Compression
+In addition to reducing search space, the proposed random freezing can also provide the benefit of reducing the model size for inference.
+Furthermore, experimental results show that the proposed method finds SLTs with better accuracy-to-model size trade-off than the SLTs obtained from dense or randomly pruned source networks.
+In particular, the SLTs found in Frozen ResNets on image classification using ImageNet significantly improve the accuracy-to-search space and accuracy-to-model size trade-offs over SLTs within dense (non-freezing) or sparse (non-locking) random networks.
+1 Introduction
+The Strong Lottery Ticket Hypothesis (SLTH) conjectured the existence of subnetworks within a randomly weighted network—strong lottery tickets (SLTs)—that achieve comparable accuracy to trained dense networks [41, 30, 25].
+The existence of such subnetworks that do not require weight training, illustrated in Figure 2(left), has been demonstrated experimentally [41, 30, 17, 21, 39] and proven theoretically for dense source networks [25, 27, 28, 6, 5, 1, 3, 8, 10].
+SLTs offer a particularly advantageous opportunity for specialized inference hardware [14] since the random weights can be reconstructed from the seed (i.e., they do not need to be memorized), and the sparse binary pruning masks—supermasks—can be greatly compressed with entropy coding, vastly reducing off-chip memory access and its associated overheads.
+Furthermore, the binary nature of both the supermasks and the random weights can be exploited for multiplier-less execution, as demonstrated practically by Hirose et al.
+[14].
+On the other hand, the existing training algorithms for finding SLTs suffer from the same high computational cost as weight optimization, since they also employ backpropagation to update one parameter per neural connection [30, 42, 6, 33].
 Figure 1: Freezing the source network by randomly pruning some parameters and locking others reduces the optimization targets for finding an SLT.
-1 Introduction
-The Strong Lottery Ticket Hypothesis (SLTH) conjectured the existence of subnetworks within a randomly weighted network—strong lottery tickets (SLTs)—that achieve comparable accuracy to trained dense networks (Zhou et al., 2019; Ramanujan et al., 2020; Malach et al., 2020).
-The existence of such subnetworks that do not require weight learning, illustrated in Figure 1(left), has been demonstrated experimentally (Zhou et al., 2019; Ramanujan et al., 2020; Huang et al., 2022; López García-Arias et al., 2023; Yeo et al., 2023) and proven theoretically for dense source networks (Malach et al., 2020; Orseau et al., 2020; Pensia et al., 2020; Diffenderfer & Kailkhura, 2021; da Cunha et al., 2021; Burkholz, 2022a; Burkholz et al., 2022; Ferbach et al., 2023; Gadhikar et al., 2023).
-Since SLTs can be reconstructed from the random seed of the weights and a supermask (a simple binary pruning mask), potentially eliminating the need for storing and loading weights from memory, they offer a particularly advantageous opportunity for specialized inference hardware (Hirose et al., 2022).
-However, the existing algorithms for finding SLTs suffer from the same high computational cost as weight optimization, since they also employ backpropagation to update one parameter per neural connection (Ramanujan et al., 2020; Zhou et al., 2021; Diffenderfer & Kailkhura, 2021; Sreenivasan et al., 2022).
+Figure 2: Freezing ($\bullet$) improves the accuracy-to-search space trade-off over pruning only (non-locking) ($\bullet$) or non-freezing ($\bullet$).
 Recently, Gadhikar et al.
-(2023) demonstrated that this steep training cost can be alleviated by randomly pruning the source network at initialization, reducing the number of edges to be optimized (see Figure 1, center).
-However, this limits the SLT search space to the high sparsity region, which may not offer the optimal accuracy.
+[10] demonstrated that this steep training cost can be alleviated by randomly pruning the source network at initialization, reducing the number of edges to be optimized (see Figure 2, center).
+However, by its nature, this method can only search SLTs in relatively high sparsity region, which also includes highly sparse networks with terrible accuracy such as layer collapse [11, 34].
+Therefore, it would be desirable if we could reduce the search space while avoiding too sparse networks.
 This paper introduces a novel method that reduces the search space by an arbitrary ratio independent of the desired SLT sparsity: in addition to random pruning at initialization, it also locks randomly chosen parameters at initialization to be a permanent part of the SLT (i.e., never pruned), as exemplified in  Figure 2(right).
 Both the pruned and the locked parameters—the frozen parameters—are excluded from the search space, therefore reducing the training cost.
-Excluding them from the supermask also reduces the model size in inference, as the random freezing pattern can also be regenerated from its seed (Wimmer et al., 2020).
-Moreover, our experiments show that SLTs found within frozen networks can achieve a better accuracy to model size trade-off than SLTs obtained from dense (Ramanujan et al., 2020; Sreenivasan et al., 2022) or sparse (Gadhikar et al., 2023) random networks.
-In addition to demonstrating this method experimentally, this paper also provides a mathematical analysis of the existence of SLTs within such partially frozen source networks.
-In summary, this paper proposes a novel method that features the following contributions:
-  • It reduces the SLT search space regardless of the desired SLT sparsity by random partial freezing (pruning and locking) of the source random network.
-  • The proposed method is theoretically supported by our SLT existence result based on subset-sum approximation with randomly frozen variables.
-  • Experimental results show that SLTs found in the frozen network achieve better accuracy-to-search space and accuracy-to-model size trade-offs than SLTs found within dense or sparse random networks.
-Figure 2: Burkholz (2022a) proved that target network of depth $L$ can be stochastically approximated by an SLT of depth $L\!+\!1$ by approximating each target weights ($w_{T}$) by a subset-sum of the source weights.
-We extend this proof to frozen source networks by solving the subset-sum approximation with some source parameters that are either unavailable ($w^{\prime}_{2}$) or must be used for the subset-sum ($w^{\prime}_{0}$).
+Far from negatively impacting performance, this cost reduction proves efficient: the SLTs found within this partial search space are more accurate than those resulting from conventional methods, as shown in Figure 2.
+Moreover, resulting from the fact that this method can also be exploited to reduce model size in inference, our experiments show that SLTs found within frozen networks can achieve a better accuracy-to-model size trade-off than SLTs obtained from dense [30, 33] or sparse [10] random networks.
+The contributions of this paper can be summarized as follows:
+  • We propose a novel method that reduces the SLT search space with any desired sparsity by random partial freezing (pruning and locking) of the source random network.
+The soundness of our method is theoretically supported by our existence theorem of SLTs in such a randomly reduced search space, based on subset-sum approximation with randomly frozen variables.
+  • We experimentally validated our method in three scenarios with highly accurate SLTs in low, medium, and high SLT sparsity regions, which reveal that parameter freezing consistently reduces the size of search spaces.
+Even with randomly frozen parameters, our method finds highly accurate SLTs that cannot be found within dense networks in some desired sparsities.
+  • Furthermore, the experimental results show that SLTs found in the frozen network achieve better accuracy-to-search space and accuracy-to-model size trade-offs than SLTs found within dense (non-freezing) or sparse (non-locking) random networks.
+As mentioned above, SLTs are quite attractive for neural engine design, as they can vastly reduce the memory size for model storage, meaning that off-chip memory access—by far the major source of energy and time consumption [15]—can be drastically reduced for energy-efficient inference acceleration [14].
+These contributions could potentially lead to attracting attention to SLTs also as an approach for efficient training, acquiring high-quality subnetworks with reduced training costs.
 2 Preliminaries
-After outlining the background on strong lottery tickets (SLTs) and their search algorithms, this section recapitulates the subset-sum approximation technique for the existence proof of SLTs (Pensia et al., 2020; da Cunha et al., 2021; Burkholz, 2022b, a) and its extension to sparse random networks (Gadhikar et al., 2023), which Section 3 expands to the existence of SLTs within frozen networks.
+After outlining the background on strong lottery tickets (SLTs) and their search algorithms, this section recapitulates the subset-sum approximation technique for the existence proof of SLTs [28, 5, 2, 1] and its extension to sparse random networks [10], which Section 3 expands to the existence of SLTs within frozen networks.
 2.1 Strong Lottery Tickets in Dense Networks
-SLTs (Zhou et al., 2019; Ramanujan et al., 2020; Malach et al., 2020) are subnetworks within a randomly weighted neural network that achieve high accuracy without any weight training.
-Compared with learned dense weight models, SLTs can be reconstructed from a small amount of information: since the random weights can be regenerated from their seed, it is only necessary to store the binary supermask and the seed (Hirose et al., 2022).
-Instead of updating weights, SLT search algorithms for deep neural networks (Ramanujan et al., 2020; Zhou et al., 2021; Sreenivasan et al., 2022) use backpropagation to update weight scores, which are then used to generate the supermask.
-For example, the Edge-Popup algorithm (Ramanujan et al., 2020; Sreenivasan et al., 2022), used in this paper, finds SLTs by applying a supermask generated by from the connections with the top-$k\%$ scores.
-The need to keep scores in addition to the weights, and the sorting of scores necessary for generating the supermask make the computational cost of Edge-Popup slightly higher than conventional weight training.
+SLTs [41, 30, 25] are subnetworks within a randomly weighted neural network that achieve high accuracy without any weight training.
+Compared with learned dense weight models, SLTs can be reconstructed from a small amount of information: since the random weights can be regenerated from their seed, it is only necessary to store the binary supermask and the seed [14].
+Instead of updating weights, SLT search algorithms for deep neural networks [30, 42, 33] use backpropagation to update weight scores, which are then used to generate the supermask.
+For example, the Edge-Popup algorithm [30, 33] used in this paper finds SLTs by applying a supermask generated from the connections with the top-$k\%$ scores.
+The need to keep scores in addition to the weights and the score sorting operation necessary for generating the supermask make the computational cost of Edge-Popup slightly higher than conventional weight training.
+Figure 3: We extend the SLT existence proof by Burkholz [1] (left) to frozen source networks (right) by solving the subset-sum approximation with some source parameters that are either unavailable ($w^{\prime}_{2}$) or must be used for the subset-sum ($w^{\prime}_{0}$).
 2.2 SLT Existence via Subset-Sum Approximation
 Given a set of random variables and a target value, the subset-sum problem consists on finding a subset whose total value approximates the target.
-Lueker (1998) showed that such subset exists with high probability if the number of random variables is sufficiently large:
-Lemma 2.1 (Subset-Sum Approximation (Lueker, 1998)).
+Lueker [24] showed that such subset exists with high probability if the number of random variables is sufficiently large:
+Lemma 2.1 (Subset-Sum Approximation [24]).
 Let $X_{1},...,X_{n}\sim U(-1,1)$ be independent, uniformly distributed random variables.
 Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by a subset-sum of $X_{i}$ if $n$ is sufficiently large.
-Based on Lemma 2.1, previous works (Pensia et al., 2020; da Cunha et al., 2021; Burkholz, 2022b, a) showed that an SLT that approximates an arbitrary target network exists in a dense source network if it is wider and deeper than the target network.
-For example, Burkholz (2022a) proved that a source network with depth $L\!+\!1$ and larger width than the target network contains an SLT that can approximate the target network with depth $L$.
+Based on Lemma 2.1, which was first introduced into the SLT context by Pensia et al.
+[28], previous works [1, 2, 5, 28] showed that an SLT that approximates an arbitrary target network exists in a dense source network if it is logarithmically wider and constantly deeper than the target network.
+In particular, Burkholz [1] proved that a source network with depth $L\!+\!1$ and larger width than the target network contains an SLT that can approximate the target network with depth $L$.
 As exemplified in Figure 3 (left), to approximate a weight $w_{T}$ of the target network, the SLT uses weights $w_{0},w_{1}$, and $w_{2}$ from the source network to construct the subset-sum.
@@
 Recently, Gadhikar et al.
-(2023) revealed that SLTs also exist within sparse source networks, i.e., random networks that have been randomly pruned at initialization (see Figure 1, center).
+[10] revealed that SLTs also exist within sparse source networks, i.e., random networks that have been randomly pruned at initialization (see Figure 2, center).
 In addition to their theoretical contributions, their approach offered a practical advantage: since there are less parameters available from the beginning, the search space is reduced in proportion to the initial pruning ratio.
-This search space reduction can be exploited for reducing the computational cost of training.
+This search space reduction can be exploited to reduce the computational cost of training.
 2.4 SLT Existence in Sparse Networks
 To prove the existence of SLTs within sparse networks, Gadhikar et al.
-(2023) extended the subset-sum approximation (Lemma 2.1) to the situation where randomly chosen variables are permanently pruned at initialization:
-Lemma 2.2 (Subset-Sum Approximation in Sparse Networks (Gadhikar et al., 2023)).
-Let $X_{1},...,X_{n}$ be as in Lemma 2.1, and $M_{1},...,M_{n}\sim Ber(p)$ be independent, Bernoulli distributed random variables with $p\in(0,1)$.Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by a subset-sum of $M_{i}X_{i}$ if $n$ is sufficiently large.
-By applying this extended lemma instead of Lemma 2.1 to the SLT existence theorem presented by Burkholz (2022a), Gadhikar et al.
-(2023) proved the SLT existence in sparse networks as the following theorem:
-Theorem 2.3 (SLT Existence in Sparse Networks (Gadhikar et al., 2023)).
-Let a target network $f_{T}$ with depth $L$ and a sparse source network $f_{S}$ with depth $L\!+\!1$.
+[10] extended the subset-sum approximation (Lemma 2.1) to the situation where randomly chosen variables are permanently pruned at initialization:
+Lemma 2.2 (Subset-Sum Approximation in Sparse Networks [10]).
+Let $X_{1},...,X_{n}$ be as in Lemma 2.1, and $M_{1},...,M_{n}\sim Ber(p)$ be independent, Bernoulli distributed random variables with $p\in(0,1)$.
+Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by a subset-sum of $M_{i}X_{i}$ if $n$ is sufficiently large.
+By applying this extended lemma instead of Lemma 2.1 to the SLT existence theorem presented by Burkholz [1], Gadhikar et al. [10] proved the SLT existence in sparse networks as the following theorem:
+Theorem 2.3 (SLT Existence in Sparse Networks [10]).
+Let a target network $f_{T}$ with depth $L$ and a sparse source network $f_{S}$ with depth $L\!+\!1$ be given.
 Assume that the source network is randomly pruned with pruning ratio $p_{l}$ for each $l$-th layer.
@@
 Although pruning at initialization provides a search space reduction, it also imposes a limitation on the sparsity range, as it can only search for SLTs with a sparsity higher than the initial pruning ratio.
-For example, although an initial pruning ratio of $90\%$ vastly reduces the search space, it limits the search to only SLTs sparser than $90\%$.
+For example, an initial pruning ratio of $90\%$ vastly reduces the search space, but it limits the search to only SLTs sparser than $90\%$.
 Therefore, this reduction in search space is incompatible with exploring the whole sparsity range for optimal accuracy.
 This paper proposes a novel method that reduces the search space regardless of the desired SLT sparsity to capitalize on the training cost reduction without sacrificing accuracy.
-We explore the existence and performance of SLTs within a frozen source network, i.e., a random network that, in addition to randomly pruned parameters, has randomly chosen parameters forced to be permanent (i.e., never pruned) part of the SLT—locked parameters (see Figure 1, right).
-Since freezing is performed at initialization, before the SLT search, frozen parameters do not have a score and are not eligible for exploration, thus reducing the search space.
-Locking allows to further reduce the parameter search without limiting the initial pruning ratio, allowing to reduce the search space even in the case of dense source networks.
-This section first provides a theoretical analysis on the existence of SLTs in such a frozen network by developing a new extension of the subset-sum approximation introduced in Section 2.
+We explore the existence and performance of SLTs within a frozen source network, i.e., a random network that (in addition to randomly pruned parameters) has randomly chosen parameters forced to be permanent part (i.e., never pruned) of the SLT—locked parameters (see Figure 2, right).
+Since freezing is performed at initialization before the SLT search, frozen parameters do not have a score and are not eligible for exploration, thus reducing the search space.
+Locking allows to further reduce the parameter search without limiting the initial pruning ratio, reducing the search space even in the case of dense source networks.
+This section first provides a theoretical analysis of the SLT existence in such frozen networks by developing a new extension of the subset-sum approximation introduced in Section 2.
 Then, based on this analysis, it proposes partial SLT search, a novel method for finding such SLTs with a reduced search space.
@@
 3.1 SLT Existence in Frozen Networks
-Here we provide a theoretical result that indicates that an SLT capable of approximating a target network exist in a frozen network.
-This is proved by extending the subset-sum approximation lemma (Lemma 2.1) to the case where some parameters are locked (for detailed proof see Appendix A).
+Here we provide a theoretical result indicating that an SLT capable of approximating a target network exists in a frozen network.
+This result is proved by extending the subset-sum approximation lemma (Lemma 2.1) to the case where some parameters are locked (for detailed proof see Appendix A).
 Lemma 3.1 (Subset-Sum Approximation in Randomly Locked Networks (Lemma A.2)).
-Let $X_{1},...,X_{n}$ be as in Lemma 2.1, and $M^{\prime}_{1},...,M^{\prime}_{n}\sim Ber(q)$ be independent, Bernoulli distributed random variables with $q\in(0,1)$.Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by the sum of $\sum_{i=1}^{n}M^{\prime}_{i}X_{i}$ and a subset-sum of $(1-M^{\prime}_{i})X_{i}$ if $n$ is sufficiently large.
-Then, combining Lemma 3.1 with Lemma 2.2 further extends the subset-sum approximation to the situation where some random variables are frozen (i.e., pruned or locked).
+Let $X_{1},...,X_{n}$ be as in Lemma 2.1, and $M^{\prime}_{1},...,M^{\prime}_{n}\sim Ber(q)$ be independent, Bernoulli distributed random variables with $q\in(0,1)$.
+Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by the sum of $\sum_{i=1}^{n}M^{\prime}_{i}X_{i}$ and a subset-sum of $(1-M^{\prime}_{i})X_{i}$ if $n$ is sufficiently large.
+Then, by combining Lemma 3.1 and Lemma 2.2, we extend the subset-sum approximation to the situation where some random variables are frozen (i.e., pruned or locked).
 Lemma 3.2 (Subset-Sum Approximation in Frozen Networks (Lemma A.3)).
@@
 Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by the sum of $\sum_{i=1}^{n}M_{i}M^{\prime}_{i}X_{i}$ and a subset-sum of $M_{i}(1-M^{\prime}_{i})X_{i}$ if $n$ is sufficiently large.
-Finally, by applying Lemma 3.2 to the Theorem 2.5 in Gadhikar et al. (2023) instead of Lemma 2.2, it follows that an SLT, which approximates a target network, exists in frozen networks.
+Finally, by applying Lemma 3.2 to the Theorem 2.5 in Gadhikar et al. [10] instead of Lemma 2.2, it follows that an SLT approximating a target network exists within a frozen network.
 Theorem 3.3 (SLT Existence in Frozen Networks (Theorem A.4)).
@@
 Figure 4: Pruning and locking set the bounds of the search space, whose optimal position is investigated.
-Figure 3: Pruning and locking set the bounds of the search space, whose optimal position is investigated.
-Figure 4: Comparison of different prune:lock proportions of a $80\%$ freezing ratio using a Conv6 on CIFAR-1010{10}10.
-Figure 5: Impact of the freezing ratio on different architectures. Pruning and locking ratios are set following Section 3.3.
 3.2 Partial SLT Search in Frozen Networks
@@
 Consequently, the pruning ratio sets the lower bound of the search space.
-On the other hand, it is not possible to search for an SLT that prunes more parameters than available, so the locking ratio sets the upper bound of the search space.
-(Specifically, the upper bound is given by $1\!-\!F_{r}$) Thus, these ratios allow to freely control the size and position of the search space over the SLT sparsity spectrum.
+On the other hand, it is not possible to search for an SLT that prunes more parameters than those available, so the locking ratio sets the upper bound of the search space.
+(Specifically, the upper bound is given by $1\!-\!L_{r}$.) Thus, these ratios allow to freely control the size and position of the search space over the SLT sparsity spectrum.
 Frozen network construction:
@@
 Our method considers the following two existing strategies for determining the layer-wise pruning ratio from the desired global pruning ratio of the source network:
-  • Edge Per Layer (EPL): Each layer’s pruning ratio is set so that they all have the same number of remaining weights (Price & Tanner, 2021; Gadhikar et al., 2023).
-  • Erdős-Rényi-Kernel (ERK): The pruning ratio of layer $l$ is proportional to the scale $(C_{in}^{(l)}+C_{out}^{(l)}+k_{h}^{(l)}+k_{w}^{(l)})/(C_{in}^{(l)}\times C_{out}^{(l)}\times k_{h}^{(l)}\times k_{w}^{(l)})$, where $C_{in}^{(l)}$, $C_{out}^{(l)}$, $k_{h}^{(l)}$, and $k_{w}^{(l)}$ denote input channels, output channels, kernel height, and kernel width of the layer $l$, respectively (Evci et al., 2020).
-The same strategy is employed to determine the layer-wise pruning and freezing ratios from their respective global ratios, and then locking ratios are calculated as the difference between the corresponding freezing and pruning ratios.
-Freezing mask compression:
+  • Edge Per Layer (EPL): Each layer’s pruning ratio is set so that they all have the same number of remaining weights [29, 10].
+  • Erdős-Rényi-Kernel (ERK): The pruning ratio of layer $l$ is proportional to the scale $(C_{in}^{(l)}+C_{out}^{(l)}+k_{h}^{(l)}+k_{w}^{(l)})/(C_{in}^{(l)}\times C_{out}^{(l)}\times k_{h}^{(l)}\times k_{w}^{(l)})$, where $C_{in}^{(l)}$, $C_{out}^{(l)}$, $k_{h}^{(l)}$, and $k_{w}^{(l)}$ denote input channels, output channels, kernel height, and kernel width of the layer $l$, respectively [7].
+The same strategy is employed to determine the layer-wise pruning and freezing ratios from their respective global ratios, and then the locking ratios are calculated as the difference between the corresponding freezing and pruning ratios.
+Freezing pattern encoding for model compression:
 The freezing pattern can be encoded during inference as a ternary mask—a freezing mask—that indicates whether a parameter is pruned, locked, or part of the supermask.
-For example, encoding pruning as $-1$, locking as $+1$, and supermask inclusion as 00, the layer-wise freezing mask can be encoded as $\bm{m}^{(l)}_{l}+(\bm{m}^{(l)}_{p}-1)$.
+For example, encoding pruning as $-1$, locking as $+1$, and supermask inclusion as 00, the layer-wise freezing mask can be encoded as $\bm{m}^{(l)}_{l}+(\bm{m}^{(l)}_{p}-1)\in\{-1,0,1\}^{n^{(l)}}$.
 Since this freezing mask is also random and fixed, it can be regenerated from its seed and ratios, similarly to the random weights.
-Furthermore, the supermask size is reduced by excluding from it the frozen parameters, so the SLTs found by this method can be compressed in inference to an even smaller memory size than those produced by the existing SLT literature (Hirose et al., 2022; Okoshi et al., 2022; López García-Arias et al., 2023).
+Furthermore, the supermask size is reduced by excluding from it the frozen parameters, so the SLTs found by this method can be compressed in inference to an even smaller memory size than those produced by the existing SLT literature [14, 26, 21].
+Also, this memory size compression reduces costly off-chip memory access on specialized neural inference accelerators [14, 4], thus SLTs within frozen networks have the potential to perform even faster and more energy-efficient inference processing.
 3.3 Optimal Positioning of the Search Space
-Here we perform a preliminary investigation of the optimal positioning for each given desired SLT sparsity by exploring different pruning:locking proportions with a fixed freezing ratio of the network.
+Figure 5: Different prune:lock proportions of a $80\%$ freezing ratio using a Conv6.
+Here we perform a preliminary investigation of the optimal positioning of the search space bounds for each given desired SLT sparsity by exploring different pruning:locking proportions with a fixed freezing ratio of the network.
 Figure 5 explores different configurations of a $80\%$ freezing ratio—a search space size of $20\%$ of the original dense source network—on a Conv6 network and compares the performance of the found SLTs on CIFAR-1010{10}10.
@@
 Consequently, hereafter the proposed method sets the global ratios in order to position the search space center as close as possible to the desired SLT sparsity, and then sets the layer-wise ratios using EPL.
-Figure 6: Freezing achieves better accuracy to search space trade-off compared to sparse ($\bullet$) or dense ($\bullet$) source networks.
-SLT sparsity set according to highest accuracy in Figure 5 with dense source networks.
 4 Experiments
-This section demonstrates that partial SLT search reduces the search space in a broad range of situations by evaluating it on image classification and node classification.
+This section demonstrates that freezing reduces the search space in a broad range of situations by evaluating it on image classification and graph node classification.
 We evaluate the impact of the freezing ratio on various network architectures, identifying three scenarios.
@@
 4.1 Experimental Settings
-The proposed partial SLT search is evaluated on image classification using the CIFAR-1010{10}10 (Krizhevsky, 2009) dataset, and on node classification using the OGBN-Arxiv (Hu et al., 2020) dataset.
-CIFAR-1010{10}10 train data is split into training and validation sets with a 4:1 ratio, while for OGBN-Arxiv we use the default set split.
-We test the models with the best validation accuracy, and report the mean of three experiment repetitions.
-For image classification we employ the VGG-like Conv6 (Simonyan & Zisserman, 2014; Ramanujan et al., 2020) and ResNet-1818{18}18 (He et al., 2016) architectures, and for graph node classification the GIN (Xu et al., 2019) architecture constructed with 4 layers modified by Huang et al.
-(2022), all implemented with no learned biases.
-ResNet-1818{18}18 and GIN use non-affine Batch Normalization (Ioffe & Szegedy, 2015), while Conv6 has no normalization layers.
-Random weights are initialized with the Kaiming Uniform distribution, while weight scores are initialized with the Kaiming Normal distribution (He et al., 2015).
-SLTs are searched using an extension of Edge-Popup (Ramanujan et al., 2020) that enforces the desired sparsity globally instead of per-layer (Sreenivasan et al., 2022).
-On CIFAR-1010{10}10, scores are optimized for $100$ epochs using stochastic gradient descent (SGD) with momentum $0.9$, batch size $128$, weight decay $0.0001$, and initial learning rates of $0.01$ and $0.1$ for Conv6 and ResNet-1818{18}18, respectively.
-On OBGN-Arxiv, scores are optimized for $400$ epochs using AdamW (Loshchilov & Hutter, 2019) with weight decay $0.0001$ and initial learning rate $0.01$.
-All experiments use cosine learning rate decay (Loshchilov & Hutter, 2017).
-Figure 7: Freezing achieves better accuracy to memory size trade-off compared to sparse ($\bullet$) or dense ($\bullet$) source networks.
+The proposed partial SLT search is evaluated on image classification using the CIFAR-1010{10}10 [19] and ImageNet [31] datasets, and on node classification using the OGBN-Arxiv [16] dataset.
+CIFAR-1010{10}10 and ImageNet train data are split into training and validation sets with a 4:1 ratio, while for OGBN-Arxiv we use the default set split.
+We test the models with the best validation accuracy and report the mean of three experiment repetitions for CIFAR-1010{10}10 and OGBN-Arxiv, and the result of one experiment for ImageNet.
+Standard deviation of experiments conducted more than once is plotted as error bars.
+For image classification we employ the VGG-like Conv6 [32, 30], ResNet [13], and Wide ResNet [40] architectures, and for graph node classification the 4-layer modified GIN [37] architecture in Huang et al.
+[17], all implemented with no learned biases.
+ResNet and GIN use non-affine Batch Normalization [18], while Conv6 has no normalization layers.
+Random weights are initialized with the Kaiming Uniform distribution, while weight scores are initialized with the Kaiming Normal distribution [12].
+SLTs are searched using an extension of Edge-Popup [30] that enforces the desired sparsity globally instead of per-layer [33].
+On CIFAR-1010{10}10, scores are optimized for $100$ epochs using stochastic gradient descent with momentum $0.9$, batch size $128$, weight decay $0.0001$, and initial learning rates of $0.01$ and $0.1$ for Conv6 and ResNet-1818{18}18, respectively.
+On ImageNet, scores are optimized by the same setting as ResNet-1818{18}18 on CIFAR-1010{10}10, but a $256$ batch size.
+On OBGN-Arxiv, scores are optimized for $400$ epochs using AdamW [23] with weight decay $0.0001$ and initial learning rate $0.01$.
+All experiments use cosine learning rate decay [22].
+These can be adequately verified with two NVIDIA H100 SXM5 94GB GPUs.
+Figure 6: Impact of the freezing ratio on different architectures. Pruning and locking ratios are set following Section 3.3.
 4.2 Reducing the Search Space with Freezing
@@
 Therefore, optimal SLTs are found in the higher sparsity range, as revealed by LABEL:fig:sparsity_acc_resnet18, benefiting of much higher pruning than locking.
-Even in this scenario, $90\%$ sparse SLTs found in frozen networks achieve similar accuracy ($<1$ point difference) while reducing the search space by as much as $80\%$.
-As an example of the scenario where optimal SLT sparsities are found in the denser range, benefitting from a higher locking ratio, we evaluate GIN in LABEL:fig:sparsity_acc_gin.
-Compared with the best-performing SLT found in the dense GIN, of $70.1\%$ accuracy and $20\%$ sparsity, by freezing $50\%$ of the search space our method finds better performing SLTs of $70.2\%$ accuracy with $40\%$ sparsity.
-These results suggest that parameter freezing at an appropriate ratio has the effect of eliminating local optimal solutions from the search space.
-It is known that Edge-Popup struggles to find the SLTs of the global optimal solution (Fischer & Burkholz, 2022).
+Even though $90\%$ of the search space is reduced in this scenario, we can find $90\%$ sparse SLTs with $88.1\%$ accuracy.
+As an example of the scenario where optimal SLT sparsities are found in the denser range, benefiting from a higher locking ratio, we evaluate GIN in LABEL:fig:sparsity_acc_gin.
+Compared with the best-performing SLT found in the dense GIN, of $70.1\%$ accuracy and $20\%$ sparsity, by freezing $50\%$ of the search space our method finds similarly performing SLTs of $69.8\%$ accuracy with $40\%$ sparsity.
+Interestingly, with low SLT sparsity (e.g., $10\%$ sparsity) in all three scenarios, despite the reduced search space, SLTs within frozen networks achieve higher accuracy than SLTs within dense networks.
+These results suggest that parameter freezing at an appropriate ratio has the effect of avoiding the inclusion of low-grade local optimal solutions in the search space.
+It is known that Edge-Popup struggles to find the SLTs of the global optimal solution [9].
 While searching for SLTs in a dense network leads to convergence to a local optimal solution, a moderate reduction of the search space may reduce the number of less accurate local optimal solutions and facilitate convergence to a more accurate local optimal solution within the reduced search space.
 In other words, we conjecture that if the network is properly frozen, the local optimal solution for the reduced search space is close to the global optimal solution of the entire search space.
-4.3 Accuracy to Search Space Trade-Off
+4.3 Accuracy-to-Search Space Trade-Off
 Section 3 sets as a requisite of the subset-sum proof that the source network has a sufficiently larger width than the target network.
@@
 SLT sparsity is fixed to that of the best performing SLT found in a dense source network in Figure 6: $50\%$ in Conv6, $90\%$ in ResNet-1818{18}18, and $20\%$ in GIN.
-Compared to the dense source network, the results show in all cases that, when the frozen network is twice as wide, the found SLTs for a given sparsity achieve similar or higher performance with a similar or smaller search space size.
-That is, the propsed method improves the accuracy to search space size trade-off.
-In the case of GIN in LABEL:fig:search_space_acc_gnn, the SLT found in a 2×2$\times$2 × wider frozen source even achieves higher accuracy than those found in any dense source network.
+The results show in all cases that the found SLTs in wide frozen networks for a given sparsity achieve similar or higher performance with a similar or smaller search space size compared to the dense source network.
+That is, the proposed method improves the accuracy-to-search space size trade-off.
 As shown in LABEL:fig:search_space_acc_conv6 and LABEL:fig:search_space_acc_gnn, in the scenarios that only allow for medium to low pruning, there is a limit to the number of parameters that can be pruned for a given SLT sparsity, and only aggressive pruning leads to less accurate SLTs.
 By additionally locking parameters, our method reduces the search space further and finds SLTs of higher accuracy than those within sparse or dense networks for a similar search space size.
-Pruning only improves the trade-off in the scenario that allows higher pruning than locking, as shown in LABEL:fig:search_space_acc_resnet18, while our method improves the accuracy to search space trade-off in all scenarios.
+Pruning only improves the trade-off in the scenario that allows higher pruning than locking, as shown in LABEL:fig:search_space_acc_resnet18, while our method improves the accuracy-to-search space trade-off in all scenarios.
+Figure 7: Freezing achieves better accuracy-to-search space trade-off compared to sparse ($\bullet$) or dense ($\bullet$) source networks.
+SLT sparsity set according to highest accuracy in Figure 6 with dense source networks.
+Figure 8: Compared to sparse ($\bullet$) or dense ($\bullet$) source networks (top-left is better), freezing achieves better accuracy-to-memory size trade-off.
 4.4 Accuracy to Model Memory Size Trade-Off
 The freezing mask compression scheme proposed in Section 3.2 allows to reduce the memory model size during inference by regenerating both the random weights and the freezing mask with a random number generator (RNG).
-Here we consider this compression and investigate the accuracy to model size trade-off offered by the SLTs found by the proposed method.
-Furthermore, considering the practical application of partial SLT search on specialized hardware, we also compare the Kaiming Uniform (KU) weight initialization used so far with the binary weights provided by the Signed Kaiming Constant (SKC) initialization (Ramanujan et al., 2020), which can be exploited for reduced computational cost in neural engines (Hirose et al., 2022).
+Here we consider this compression and investigate the accuracy-to-model size trade-off offered by the SLTs found by the proposed method.
+Furthermore, we also compare the Kaiming Uniform (KU) weight initialization used so far with the binary weights provided by the Signed Kaiming Constant (SKC) initialization [30], which can be exploited for reduced computational cost in neural engines [14].
 For SKC we scale weights by $1/\sqrt{1-k_{l}}$, where $k_{l}$ is the sparsity of each layer, as proposed by Ramanujan et al.
-(2020).
-As illustrated in Figure 7, compared to the SLTs found in dense or sparse source networks, SLTs found by our method achieve similar or higher accuracy for similar or smaller model size, thus improving the accuracy to model size trade-off in all scenarios.
+[30].
+As illustrated in Figure 8, compared to the SLTs found in dense or sparse source networks, SLTs found by our method achieve similar or higher accuracy for similar or smaller model size, thus improving the accuracy-to-model size trade-off in all scenarios.
 In other words, partial SLT search not only reduces the search space during training, but also reduces the model size during inference, further improving performance over conventional SLT search methods.
-Empirically, it is known that SLTs within a SKC initialized dense network achieve better performance than with continuous random weights (Ramanujan et al., 2020; Okoshi et al., 2022; López García-Arias et al., 2023; Yan et al., 2023).
+Empirically, it is known that SLTs within a SKC initialized dense network achieve better performance than with continuous random weights [30, 26, 21, 38].
 Our results show that such a trend can also be observed in frozen source networks.
 Nonetheless, in all source networks, we find that this improvement is smaller the wider the source network is, suggesting that the requirement for source networks of larger width is weaker in the case of binary weights.
-Table 1: Comparison between trained-weight networks and SLTs found in dense, sparse, and frozen networks.
-Conv6 & CIFAR-1010{10}10
-Method
-(Source Network)
-Weight
-Initialization
-Sparsity
-Pruning
-Ratio [%]
-Locking
-Ratio [%]
-Top-1 Test
-Acc. [%]
-Search Space
-[# Parameters]
-Model Size
-[MB]
-Weight Training KU - - - $87.1$ $2.26$ M $8.63$
-SLT (Dense) SKC $50$ 00 00 $86.2$ $2.26$ M $0.27$
-SLT (Sparse) SKC $50$ $45$ 00 $66.7$ $1.24$ M $0.15$
-SLT (Frozen) SKC $50$ $25$ $25$ $84.8$ $1.13$ M $0.13$
-ResNet-${18}$ CIFAR-1010{10}10
-Weight Training KU - - - $92.5$ $11.16$ M $42.63$
-SLT (Dense) SKC $90$ 00 00 $91.0$ $11.16$ M $1.37$
-SLT (Sparse) SKC $90$ $85$ 00 $89.9$ $1.67$ M $0.24$
-SLT (Frozen) SKC $90$ $82.5$ $2.5$ $89.8$ $1.67$ M $0.24$
-SLT (Frozen) SKC $90$ $85$ $5$ $88.9$ $1.12$ M $0.17$
-GIN & OGBN-Arxiv
-Weight Training KU - - - $70.1$ $0.38$ M $1.452$
-SLT (Dense) SKC $20$ 00 00 $70.1$ $0.38$ M $0.054$
-SLT (Sparse) SKC $20$ $15$ 00 $63.1$ $0.32$ M $0.047$
-SLT (Frozen) SKC $20$ 00 $40$ $70.6$ $0.23$ M $0.036$
-4.5 Result Analysis
-Table 1 summarizes the presented results and compares the proposed method to weight training and SLT search in dense and sparse source networks for the same network architecture and SLT sparsity.
-In the case of Conv6, compared to weight training our method provides reductions of 2×2$\times$2 × and 66.4×66.4$\times$66.4 × of the search space and model size, respectively, in exchange of a small accuracy drop.
-Compared to the SLT found in the sparse source network, the SLT found by partial search achieves much higher accuracy.
-[…]
-Interestingly, the SLT found in the frozen GIN model achieves higher accuracy than the trained-weight network, even though the model size and search space are reduced by 40.3×40.3$\times$40.3 × and 1.7×1.7$\times$1.7 ×, respectively.
-The SLT found within a dense network also achieves comparable accuracy to the trained-weight network, but the SLT found by our method performs better and is smaller.
-Generally, graph neural networks (GNNs) suffer from a generalization performance degradation due to the over-smoothing problem (Li et al., 2018).
-Searching for SLTs within dense GNNs has been shown to mitigate the over-smoothing problem (Huang et al., 2022) and achieve higher accuracy.
-As our method is even more accurate than dense networks, it is possible that parameter freezing offers an even stronger solution to the over-smoothing problem.
+4.5 ImageNet Experiments
+Figure 9: Freezing enhances the accuracy-to-model size trade-off even on a practical setting with large-scale dataset and models.
+Finally, we evaluate our method using larger models on a large-scale dataset: deeper and wider ResNets on ImageNet.
+Since SLTs with $80\%$ sparsity achieve the highest accuracy in a dense source ResNet-5050{50}50 (for details, see Figure 10 in Appendix B), we compare the three methods using $80\%$ SLT sparsity.
+Figures 2 and 9 compare the accuracy-to-search space and accuracy-to-model size trade-offs, respectively, in finding the $80\%$ sparsity SLTs between the proposed and conventional methods with SKC using ImageNet.
+In both of the figures, such SLTs searched from non-frozen (blue), $70\%$ pruned (orange), and $65\%$ pruned and $5\%$ locked ($70\%$ frozen, green) are compared.
+Despite the more challenging setting and the significant $70\%$ search space reduction, our method finds SLTs that are substantially more accurate than other methods for the same search space size (Figure 2) and for the same model memory size (Figure 9).
+This result demonstrates that the effective combination of parameter pruning and locking at initialization can improve the SLT search efficiency even on large-scale datasets and models.
+Notably, Figure 9 reveals that the proposed method enhances the accuracy-to-model size trade-off, validating our claim it could be deployed on specialized inference hardware, such as the one like Hirose et al.
+[14], to enhance the efficiency of practical SLT applications.
+5 Limitations
+Our results have the following limitations: 1) Although the proposed method can effectively reduce the search space for training and the memory model size for inference in principle, actual hardware implementation remains for future work.
+The previous work of the SLT-specialized neural engine, Hiddenite [14], would be a promising direction for such future implementation.
+2) This paper provides theoretical support for the existence of SLTs in frozen networks, but it does not provide their superiority in the approximation capability compared to random pruning, a special case of random freezing, which may depend on the sparsity nature of target networks.
+Also, the theoretical consideration of the appropriate proportion of these ratios is left for future work.
+3) Even though we demonstrate our method on various model architectures including CNNs, ResNet families and GIN, following previous SLT work [41, 30, 21, 17, 38], we have to leave it for future work to apply our method to Transformers [35] because SLT itself has not yet been established for Transformers.
 6 Conclusion
-This paper capitalizes on the fact that pruning a randomly weighted network at initialization reduces the SLT search space, but identifies that it limits the search to a suboptimal sparsity region.
+This paper capitalizes on the fact that pruning a randomly weighted network at initialization reduces the SLT search space, but identifies that doing so limits the search to a suboptimal sparsity region.
 This problem is tackled by freezing (i.e., pruning or locking) some parameters at initialization, excluding them from the search.
-Freezing allows to search for SLTs efficiently in the optimal sparsity range, while further reducing search space during training and model size during inference.
-Experimental results show that SLTs found in frozen networks improve the accuracy to search space and accuracy to model size trade-offs compared to SLTs found in dense (Ramanujan et al., 2020; Sreenivasan et al., 2022) or sparse source networks (Gadhikar et al., 2023).
-Interestingly, although for weight training random parameter locking has only been found useful for reducing accuracy degradation in network compression (Zhou et al., 2019; Wimmer et al., 2020), we identify a scenario in SLT training where it can be used for raising accuracy, even outperforming weight-trained networks.
-Our method can be interpreted as being capable of generating more useful SLT information from a random seed than previous methods, and as a result, it offers an opportunity for reducing off-chip memory access in specialized SLT accelerators (Hirose et al., 2022).
-We hope that this work will serve as a base for raising further the efficiency of SLT training and inference.
-Impact Statements
-This paper presents work whose goal is to advance the field of machine learning.
-Among the many potential societal consequences of our work, we highlight its potential to reduce the computational cost of inference and training of neural networks, and thus their energy consumption footprint.
+Freezing allows to search for SLTs efficiently in the optimal sparsity range, while further reducing the search space during training and the model size during inference.
+Experimental results show that SLTs found in frozen networks improve the accuracy-to-search space and accuracy-to-model size trade-offs compared to SLTs found in dense [30, 33] or sparse source networks [10].
+Interestingly, although for weight training random parameter locking has only been found useful for reducing accuracy degradation in network compression [41, 36], we identify scenarios in SLT training where they can be used for raising accuracy.
+Our method can be interpreted as being capable of generating more useful SLT information from a random seed than previous methods, offering an opportunity for reducing off-chip memory access in specialized SLT accelerators [14].
+We hope this work will serve as a base for raising the efficiency of SLT training and inference.
+Acknowledgments and Disclosure of Funding
+This work was supported in part by JSPS KAKENHI Grant Number JP23H05489.
 Appendix A Proof for Strong Lottery Tickets (SLTs) Existence in Frozen Networks
-This section describes the proofs of the lemma and theorem introduced in the manuscript. For reference, we present a theorem presented by Gadhikar et al. (2023) in advance.
-Theorem A.1 (SLT Existence in Sparse Networks (Gadhikar et al., 2023)).
+This section describes the proofs of the lemma and theorem introduced in the manuscript. For reference, we present a theorem presented by Gadhikar et al. [10] in advance.
+Theorem A.1 (SLT Existence in Sparse Networks [10]).
 Let input data $\mathcal{D}$ , a target network $f_{T}$ with depth $L$, and a source network $f_{S}$ with depth $L+1$, parameters $\bm{\theta}_{S}$, and edge probabilities $p_{l}$ be given.
@@
 where $C>0$ denotes a generic constant that is independent of $n_{T,l},L,p_{l},\delta$, and $\epsilon$.
-We first extend the lemma of the subset-sum approximation shown by Lueker (1998) to the case where some random variables are always included in the subset-sum.
+We first extend the lemma of the subset-sum approximation shown by Lueker [24] to the case where some random variables are always included in the subset-sum.
 Lemma A.2 (Subset-Sum Approximation in Randomly Locked Networks).
@@
 where the former part is used for approximating $z\in[-1,1]$ and the latter part for approximating $\sum_{i=1}^{n}M_{i}X_{i}$.
-To approximate $z\in[-1,1]$, we can directly apply Corollary 2.5 from Lueker (1998):
+To approximate $z\in[-1,1]$, we can directly apply Corollary 2.5 from Lueker [24]:
 $\mathbb{P}\left(\forall z\in\left[-1,1\right],\exists I\subset\{1,\cdots,n_{1}\}\text{ s.t. }|z-\sum_{i\in I}X^{\prime}_{i}|\geq\varepsilon\right)\geq 1-\delta$ (11)
@@
 Since $m\leq\beta n$ holds, $n\geq\frac{2\log(2/\delta)}{3\alpha^{2}\beta}$ is enough.
-From the proof of Corollary 3.1 in Lueker (1998), for any $\gamma\in(0,\frac{1}{4})$, we know that
+From the proof of Corollary 3.1 in Lueker [24], for any $\gamma\in(0,\frac{1}{4})$, we know that
 $\mathbb{P}\left(\forall z\in\left[-\gamma n_{2},\gamma n_{2}\right],\exists I\subset\{1,\cdots,n_{2}\}\text{ s.t. }|z-\sum_{i\in I}X^{\prime\prime}_{i}|\geq\varepsilon\right)$ (13)
@@
 By combining (11), (12), (15) with the fact that $n=m+n_{1}+n_{2}$, and replacing $\delta$ with $\delta/4$, we obtain the desired results. ∎
-By similar arguments as in Lueker (1998) or Pensia et al.
-(2020), we can easily generalize Lemma A.2 to the case where the distribution followed by $X_{i}$ contains the uniform distribution.
+By similar arguments as in Lueker [24] or Pensia et al.
+[28], we can easily generalize Lemma A.2 to the case where the distribution followed by $X_{i}$ contains the uniform distribution.
 Also, by combining our proof and the proof of Lemma 2.4 in Gadhikar et al.
-(2023), we can prove the following extension of Lemma A.2:
+[10], we can prove the following extension of Lemma A.2:
 Lemma A.3 (Subset-Sum Approximation in Frozen Networks.).
@@
 Proof.
-By replacing Lemma 2.4 in Gadhikar et al. (2023) with our Lemma A.3, we obtain the desired result as an extension of Theorem 2.5 in Gadhikar et al. (2023). ∎
+By replacing Lemma 2.4 in Gadhikar et al. [10] with our Lemma A.3, we obtain the desired result as an extension of Theorem 2.5 in Gadhikar et al. [10]. ∎
+Appendix B Additional Experimental Results
+B.1 Accuracy of SLT within Dense ResNet-50
+This section introduces the preliminary experiment used for determining the main experimental setup.
+Figure 10 compares the accuracy of SLTs within a dense ResNet-5050{50}50 source at different sparsity using ImageNet.
+The SLT with $80\%$ sparsity is the most accurate, reaching $66.8\%$ accuracy.
+Figure 10: Accuracy comparison of SLTs of different sparsity within a dense ResNet-5050{50}50 on ImageNet.
+B.2 Result Analysis
+Table 1 summarizes the presented results and compares the proposed method to weight training and SLT search in dense and sparse source networks for the same network architecture and SLT sparsity.
+In the case of Conv6, compared to weight training our method provides reductions of 2×2$\times$2 × and 66.4×66.4$\times$66.4 × of the search space and model size, respectively, in exchange of a small accuracy drop.
+Compared to the SLT found in the sparse source network, the SLT found by partial search achieves much higher accuracy.
+[…]
+Interestingly, the SLT found in the frozen GIN model achieves comparable accuracy to the trained-weight network, even though the model size and search space are reduced by 40.3×40.3$\times$40.3 × and 1.7×1.7$\times$1.7 ×, respectively.
+The SLT found within a dense network also achieves comparable accuracy to the trained-weight network, but the SLT found by our method is smaller.
+Generally, graph neural networks (GNNs) suffer from a generalization performance degradation due to the over-smoothing problem [20].
+Searching for SLTs within dense GNNs has been shown to mitigate the over-smoothing problem [17] and achieve higher accuracy.
+As our method is similarly accurate compared to dense networks, it is possible that parameter freezing offers an even stronger solution to the over-smoothing problem.
+Although the ImageNet experiment falls into the scenario where a relatively sparse SLT is more accurate (like the ResNet-1818{18}18 experiment on CIFAR-1010{10}10), the SLT found in a sparse network is significantly less accurate.
+On the other hand, the inclusion of parameter locking shows a significant improvement in SLT accuracy.
+These results suggest that the number of highly accurate SLT patterns decreases as the difficulty of the problem increases, and the way the strategy of search space reduction affects the performance of the found SLTs more severely.
+As also shown in Figure 2 in the main manuscript, the SLT found in the frozen Wide ResNet-5050{50}50 achieves $1.4$ points higher accuracy than the SLT found in the dense ResNet-3434{34}34, even though the search space is smaller, indicating superiority of our method in improving the accuracy-to-search space trade-off.
+Table 1: Comparison between trained-weight networks and SLTs found in dense, sparse, and frozen networks.
+Conv6 & CIFAR-1010{10}10
+Method
+(Source Net.)
+Weight
+Init.
+Sparsity
+Pruning
+Ratio [%]
+Locking
+Ratio [%]
+Top-1 Test
+Acc. [%]
+Search Space
+[# Param.]
+Model
+Size [MB]
+Weight Training KU - - - $87.1$ $2.26$ M $8.63$
+SLT (Dense) SKC $50$ 00 00 $86.2$ $2.26$ M $0.27$
+SLT (Sparse) SKC $50$ $45$ 00 $66.7$ $1.24$ M $0.15$
+SLT (Frozen) SKC $50$ $25$ $25$ $84.8$ $1.13$ M $0.13$
+ResNet-${18}$ CIFAR-1010{10}10
+Weight Training KU - - - $92.4$ $11.16$ M $42.63$
+SLT (Dense) SKC $90$ 00 00 $91.0$ $11.16$ M $1.37$
+SLT (Sparse) SKC $90$ $85$ 00 $89.9$ $1.67$ M $0.24$
+SLT (Frozen) SKC $90$ $82.5$ $2.5$ $89.8$ $1.67$ M $0.24$
+SLT (Frozen) SKC $90$ $85$ $5$ $88.9$ $1.12$ M $0.17$
+GIN & OGBN-Arxiv
+Weight Training KU - - - $70.1$ $0.38$ M $1.452$
+SLT (Dense) SKC $20$ 00 00 $70.0$ $0.38$ M $0.054$
+SLT (Sparse) SKC $20$ $15$ 00 $62.3$ $0.32$ M $0.047$
+SLT (Frozen) SKC $20$ 00 $40$ $69.2$ $0.23$ M $0.036$
+ResNet-${50}$ ImageNet
+Weight Training KU - - - $74.4$ $25.50$ M $97.49$
+SLT (Dense) SKC $80$ 00 00 $66.8$ $25.50$ M $3.24$
+SLT (Sparse) SKC $80$ $70$ 00 $47.1$ $7.65$ M $1.11$
+SLT (Frozen) SKC $80$ $65$ $5$ $55.2$ $7.65$ M $1.11$
+ResNet-${34}$ ImageNet
+SLT (Dense) SKC $80$ 00 00 $62.5$ $21.78$ M $2.66$
+SLT (Sparse) SKC $80$ $70$ 00 $40.7$ $6.53$ M $0.84$
+SLT (Frozen) SKC $80$ $65$ $5$ $51.2$ $6.53$ M $0.84$
+ResNet-${18}$ ImageNet
+SLT (Dense) SKC $80$ 00 00 $54.0$ $11.68$ M $1.43$
+SLT (Sparse) SKC $80$ $70$ 00 $25.2$ $3.50$ M $0.45$
+SLT (Frozen) SKC $80$ $65$ $5$ $39.4$ $3.50$ M $0.45$
+Wide ResNet-${50}$ ImageNet
+SLT (Dense) SKC $80$ 00 00 $70.8$ $68.81$ M $8.46$
+SLT (Sparse) SKC $80$ $70$ 00 $57.7$ $20.64$ M $2.72$
+SLT (Frozen) SKC $80$ $65$ $5$ $63.9$ $20.64$ M $2.72$
+Appendix C Impact Statements
+This paper presents work whose goal is to advance the field of machine learning.
+Among the many potential societal consequences of our work, we highlight its potential to reduce the computational cost of inference and training of neural networks, and thus their energy consumption footprint.

### v3

@@
-Partial Search in a Frozen Network is Enough to Find a Strong Lottery Ticket
+Partially Frozen Random Networks Contain Compact Strong Lottery Tickets
 Abstract
@@
 Recently, Gadhikar et al.
-[10] demonstrated that SLTs can also be found within a randomly pruned source network, thus reducing the SLT search space.
-However, this limits the search to SLTs that are even sparser than the source, leading to worse accuracy due to unintentionally high sparsity.
-This paper proposes a method that reduces the SLT search space by an arbitrary ratio independent of the desired SLT sparsity.
-A random subset of the initial weights is excluded from the search space by freezing it—i.e., by either permanently pruning them or locking them as a fixed part of the SLT.
-In addition to reducing search space, the proposed random freezing can also provide the benefit of reducing the model size for inference.
-Furthermore, experimental results show that the proposed method finds SLTs with better accuracy-to-model size trade-off than the SLTs obtained from dense or randomly pruned source networks.
-In particular, the SLTs found in Frozen ResNets on image classification using ImageNet significantly improve the accuracy-to-search space and accuracy-to-model size trade-offs over SLTs within dense (non-freezing) or sparse (non-locking) random networks.
+(2023) demonstrated that SLTs could also be found within a randomly pruned source network.
+This phenomenon can be exploited to further compress the small memory size required by SLTs.
+However, their method is limited to SLTs that are even sparser than the source, leading to worse accuracy due to unintentionally high sparsity.
+This paper proposes a method for reducing the SLT memory size without restricting the sparsity of the SLTs that can be found.
+A random subset of the initial weights is frozen by either permanently pruning them or locking them as a fixed part of the SLT, resulting in a smaller model size.
+Experimental results show that Edge-Popup (Ramanujan et al., 2020; Sreenivasan et al., 2022) finds SLTs with better accuracy-to-model size trade-off within frozen networks than within dense or randomly pruned source networks.
+In particular, freezing $70\%$ of a ResNet on ImageNet provides $\times3$ compression compared to the SLT found within a dense counterpart, raises accuracy by up to $14.12$ points compared to the SLT found within a randomly pruned counterpart, and offers a better accuracy-model size trade-off than both.
 1 Introduction
-The Strong Lottery Ticket Hypothesis (SLTH) conjectured the existence of subnetworks within a randomly weighted network—strong lottery tickets (SLTs)—that achieve comparable accuracy to trained dense networks [41, 30, 25].
-The existence of such subnetworks that do not require weight training, illustrated in Figure 2(left), has been demonstrated experimentally [41, 30, 17, 21, 39] and proven theoretically for dense source networks [25, 27, 28, 6, 5, 1, 3, 8, 10].
-SLTs offer a particularly advantageous opportunity for specialized inference hardware [14] since the random weights can be reconstructed from the seed (i.e., they do not need to be memorized), and the sparse binary pruning masks—supermasks—can be greatly compressed with entropy coding, vastly reducing off-chip memory access and its associated overheads.
-Furthermore, the binary nature of both the supermasks and the random weights can be exploited for multiplier-less execution, as demonstrated practically by Hirose et al.
-[14].
-On the other hand, the existing training algorithms for finding SLTs suffer from the same high computational cost as weight optimization, since they also employ backpropagation to update one parameter per neural connection [30, 42, 6, 33].
-Figure 1: Freezing the source network by randomly pruning some parameters and locking others reduces the optimization targets for finding an SLT.
-Figure 2: Freezing ($\bullet$) improves the accuracy-to-search space trade-off over pruning only (non-locking) ($\bullet$) or non-freezing ($\bullet$).
+The strong lottery ticket hypothesis (SLTH) conjectured the existence of subnetworks within a randomly weighted network—strong lottery tickets (SLTs)—that achieve comparable accuracy to trained dense networks (Zhou et al., 2019; Ramanujan et al., 2020; Malach et al., 2020).
+The existence of such subnetworks that do not require weight training, illustrated in Figure 2 (left), has been demonstrated experimentally (Zhou et al., 2019; Ramanujan et al., 2020; Huang et al., 2022; López García-Arias et al., 2023; Yeo et al., 2023) and proven theoretically for dense source networks (Malach et al., 2020; Orseau et al., 2020; Pensia et al., 2020; Diffenderfer & Kailkhura, 2021; da Cunha et al., 2021; Burkholz, 2022a; Burkholz et al., 2022; Ferbach et al., 2023; Gadhikar et al., 2023).
+They appear in networks with excessive amounts of parameters, and their existence indicates the possibility of new optimization strategies for deep neural networks based on a network connectivity perspective.
+SLTs offer a particularly advantageous opportunity for specialized inference hardware (Hirose et al., 2022; Chen et al., 2022; 2023) since the random weights can be reconstructed from the seed (i.e., they do not need to be memorized), and the binary mask—supermask—can be greatly compressed with entropy coding, vastly reducing off-chip memory access and its associated overheads.
+Furthermore, the binary nature of both the supermask and the random weights can be exploited for multiplier-less execution, as demonstrated practically by Hirose et al.
+(2022).
+Figure 1: Freezing the source network by randomly pruning some parameters and locking others reduces the memorized supermask for finding an SLT.
+Figure 2: Freezing ($\bullet$) improves the accuracy-to-model size trade-off over pre-pruning only ($\bullet$) or non-freezing ($\bullet$).
 Recently, Gadhikar et al.
-[10] demonstrated that this steep training cost can be alleviated by randomly pruning the source network at initialization, reducing the number of edges to be optimized (see Figure 2, center).
-However, by its nature, this method can only search SLTs in relatively high sparsity region, which also includes highly sparse networks with terrible accuracy such as layer collapse [11, 34].
-Therefore, it would be desirable if we could reduce the search space while avoiding too sparse networks.
-This paper introduces a novel method that reduces the search space by an arbitrary ratio independent of the desired SLT sparsity: in addition to random pruning at initialization, it also locks randomly chosen parameters at initialization to be a permanent part of the SLT (i.e., never pruned), as exemplified in  Figure 2(right).
-Both the pruned and the locked parameters—the frozen parameters—are excluded from the search space, therefore reducing the training cost.
-Far from negatively impacting performance, this cost reduction proves efficient: the SLTs found within this partial search space are more accurate than those resulting from conventional methods, as shown in Figure 2.
-Moreover, resulting from the fact that this method can also be exploited to reduce model size in inference, our experiments show that SLTs found within frozen networks can achieve a better accuracy-to-model size trade-off than SLTs obtained from dense [30, 33] or sparse [10] random networks.
-The contributions of this paper can be summarized as follows:
-  • We propose a novel method that reduces the SLT search space with any desired sparsity by random partial freezing (pruning and locking) of the source random network.
-The soundness of our method is theoretically supported by our existence theorem of SLTs in such a randomly reduced search space, based on subset-sum approximation with randomly frozen variables.
-  • We experimentally validated our method in three scenarios with highly accurate SLTs in low, medium, and high SLT sparsity regions, which reveal that parameter freezing consistently reduces the size of search spaces.
-Even with randomly frozen parameters, our method finds highly accurate SLTs that cannot be found within dense networks in some desired sparsities.
-  • Furthermore, the experimental results show that SLTs found in the frozen network achieve better accuracy-to-search space and accuracy-to-model size trade-offs than SLTs found within dense (non-freezing) or sparse (non-locking) random networks.
-As mentioned above, SLTs are quite attractive for neural engine design, as they can vastly reduce the memory size for model storage, meaning that off-chip memory access—by far the major source of energy and time consumption [15]—can be drastically reduced for energy-efficient inference acceleration [14].
-These contributions could potentially lead to attracting attention to SLTs also as an approach for efficient training, acquiring high-quality subnetworks with reduced training costs.
+(2023) showed that accurate SLTs could be found even if the number of edges to be optimized was reduced by randomly pruning the source network at initialization (see Figure 2, center).
+Since the random pre-pruning mask can be reconstructed from the seed in the same way as the random weights, and thus there is no need to store the part of the supermask corresponding to pre-pruned weights, their approach can be exploited to further reduce the memory cost required by the SLT in specialized hardware.
+However, when aiming for high compression, their method can only search for SLTs in the relatively high sparsity region, and may even lead to layer collapse (Hayou et al., 2020; Tanaka et al., 2020).
+Furthermore, since the SLT sparsity regions where highly accurate SLTs exist depend on the dataset and network architecture, a search limited to sparse regions may fail to find accurate SLTs.
+For instance, work on SLTs within graph neural networks has shown that dense SLTs are more accurate than sparse ones in some settings (Huang et al., 2022; Yan et al., 2023).
+Therefore, it would be desirable to use a method that allows for increased randomness in SLTs for further model compression, but has the freedom to allocate it in the SLT sparsity regions that lead to more accurate tickets.
+This paper introduces such a novel method to reduce the memory cost of the optimized supermask without restricting the desired sparsity of SLTs to be searched for: in addition to random pruning at initialization, it also locks randomly chosen parameters at initialization to be a permanent part of the SLT (i.e., never pruned), as exemplified in Figure 2 (right).
+Both the randomly pruned and the locked parameters—the frozen parameters—are left completely random and can be regenerated from seed.
+The weights corresponding to the optimized supermask region (less than $50\%$ of the total) are also reconstructed from seed, and the supermask is binary and sparse, producing a highly compressible model.
+Far from negatively impacting performance, this cost reduction is efficient: as shown in Figure 2, the frozen SLTs achieve a higher accuracy than SLTs with a similar size resulting from conventional methods.
+The contributions of this paper are summarized as follows:
+  • We propose a novel method that vastly reduces the number of parameters to be memorized for finding an accurate SLT by freezing (pruning and locking) the source random network.
+  • We experimentally validate our method in three scenarios corresponding to low, medium, and high optimal SLT sparsity regions, which reveal that parameter freezing consistently produces smaller yet accurate supermasks.
+Even with randomly frozen parameters, we find highly accurate SLTs that cannot be found within dense networks for some desired sparsities.
+  • Furthermore, the experimental results show that SLTs found in frozen networks achieve comparable or better accuracy-to-model size trade-off than SLTs found within dense (non-freezing) or sparse (non-locking) random networks.
+As mentioned above, SLTs are quite attractive for neural engine design, as they can vastly reduce the memory size for model storage, meaning that off-chip memory access—by far the major bottleneck of energy and time consumption (Horowitz, 2014)—can be drastically reduced for energy-efficient inference acceleration (Hirose et al., 2022).
+Our contributions have the potential to reduce off-chip memory access further and to make inference more energy-efficient than previous designs.
 2 Preliminaries
-After outlining the background on strong lottery tickets (SLTs) and their search algorithms, this section recapitulates the subset-sum approximation technique for the existence proof of SLTs [28, 5, 2, 1] and its extension to sparse random networks [10], which Section 3 expands to the existence of SLTs within frozen networks.
+This section outlines the background of strong lottery tickets (SLTs) within dense or sparse networks and SLT search algorithms.
 2.1 Strong Lottery Tickets in Dense Networks
-SLTs [41, 30, 25] are subnetworks within a randomly weighted neural network that achieve high accuracy without any weight training.
-Compared with learned dense weight models, SLTs can be reconstructed from a small amount of information: since the random weights can be regenerated from their seed, it is only necessary to store the binary supermask and the seed [14].
-Instead of updating weights, SLT search algorithms for deep neural networks [30, 42, 33] use backpropagation to update weight scores, which are then used to generate the supermask.
-For example, the Edge-Popup algorithm [30, 33] used in this paper finds SLTs by applying a supermask generated from the connections with the top-$k\%$ scores.
-The need to keep scores in addition to the weights and the score sorting operation necessary for generating the supermask make the computational cost of Edge-Popup slightly higher than conventional weight training.
-Figure 3: We extend the SLT existence proof by Burkholz [1] (left) to frozen source networks (right) by solving the subset-sum approximation with some source parameters that are either unavailable ($w^{\prime}_{2}$) or must be used for the subset-sum ($w^{\prime}_{0}$).
-2.2 SLT Existence via Subset-Sum Approximation
-Given a set of random variables and a target value, the subset-sum problem consists on finding a subset whose total value approximates the target.
-Lueker [24] showed that such subset exists with high probability if the number of random variables is sufficiently large:
-Lemma 2.1 (Subset-Sum Approximation [24]).
+SLTs (Zhou et al., 2019; Ramanujan et al., 2020; Malach et al., 2020) are subnetworks within a randomly weighted neural network that achieve high accuracy without any weight training.
+Compared with learned dense weight models, SLTs can be reconstructed from a small amount of information: since the random weights can be regenerated from their seed, it is only necessary to store the binary supermask and the seed (Hirose et al., 2022).
+SLT search algorithms for deep neural networks (Zhou et al., 2019; Ramanujan et al., 2020; Zhou et al., 2021; Sreenivasan et al., 2022) find SLTs by updating weight scores, which are then used to generate the supermask, instead of updating weights.
+For example, the Edge-Popup algorithm (Ramanujan et al., 2020; Sreenivasan et al., 2022) used in this paper finds SLTs by applying a supermask generated from the connections with the top-$k\%$ scores.
+SLT Existence via Subset-Sum Approximation
+Given a set of random variables and a target value, the subset-sum problem consists of finding a subset whose total value approximates the target.
+Lueker (1998) showed that such a subset exists with high probability if the number of random variables is sufficiently large:
+Lemma 2.1 (Subset-Sum Approximation (Lueker, 1998)).
 Let $X_{1},...,X_{n}\sim U(-1,1)$ be independent, uniformly distributed random variables.
 Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by a subset-sum of $X_{i}$ if $n$ is sufficiently large.
-Based on Lemma 2.1, which was first introduced into the SLT context by Pensia et al.
-[28], previous works [1, 2, 5, 28] showed that an SLT that approximates an arbitrary target network exists in a dense source network if it is logarithmically wider and constantly deeper than the target network.
-In particular, Burkholz [1] proved that a source network with depth $L\!+\!1$ and larger width than the target network contains an SLT that can approximate the target network with depth $L$.
-As exemplified in Figure 3 (left), to approximate a weight $w_{T}$ of the target network, the SLT uses weights $w_{0},w_{1}$, and $w_{2}$ from the source network to construct the subset-sum.
-Therefore, the source network is required to be sufficiently wider than the target network, as stated in Lemma 2.1.
+Based on this Lemma 2.1, which was first introduced into the SLT context by Pensia et al.
+(2020), previous works (Burkholz, 2022a; b; da Cunha et al., 2021; Pensia et al., 2020) showed that an SLT that approximates an arbitrary target network exists in a dense source network if it is logarithmically wider and constantly deeper than the target network.
+In particular, Burkholz (2022a) proved that a source network with depth $L\!+\!1$ and larger width than the target network contains an SLT that can approximate the target network with depth $L$.
 2.2 Strong Lottery Tickets in Sparse Networks
 Recently, Gadhikar et al.
-[10] revealed that SLTs also exist within sparse source networks, i.e., random networks that have been randomly pruned at initialization (see Figure 2, center).
-In addition to their theoretical contributions, their approach offered a practical advantage: since there are less parameters available from the beginning, the search space is reduced in proportion to the initial pruning ratio.
-This search space reduction can be exploited to reduce the computational cost of training.
-2.4 SLT Existence in Sparse Networks
+(2023) revealed that SLTs also exist within sparse source networks, i.e., random networks that have been randomly pruned at initialization (see Figure 2, center).
+They showed its existence experimentally as well as theoretically, as outlined later: an SLT that approximates a given target network exists with high probability in a sparse source network that is sufficiently wider and deeper than the target.
+SLT Existence in Sparse Networks
 To prove the existence of SLTs within sparse networks, Gadhikar et al.
-[10] extended the subset-sum approximation (Lemma 2.1) to the situation where randomly chosen variables are permanently pruned at initialization:
-Lemma 2.2 (Subset-Sum Approximation in Sparse Networks [10]).
+(2023) extended the subset-sum approximation (Lemma 2.1) to the situation where randomly chosen variables are permanently pruned at initialization:
+Lemma 2.2 (Subset-Sum Approximation in Sparse Networks (Gadhikar et al., 2023)).
 Let $X_{1},...,X_{n}$ be as in Lemma 2.1, and $M_{1},...,M_{n}\sim Ber(p)$ be independent, Bernoulli distributed random variables with $p\in(0,1)$.
 Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by a subset-sum of $M_{i}X_{i}$ if $n$ is sufficiently large.
-By applying this extended lemma instead of Lemma 2.1 to the SLT existence theorem presented by Burkholz [1], Gadhikar et al. [10] proved the SLT existence in sparse networks as the following theorem:
-Theorem 2.3 (SLT Existence in Sparse Networks [10]).
+By applying this extended lemma instead of Lemma 2.1 to the SLT existence theorem presented by Burkholz (2022a), Gadhikar et al.
+(2023) proved the SLT existence in sparse networks as following theorem:
+Theorem 2.1 (SLT Existence in Sparse Networks (Gadhikar et al., 2023)).
 Let a target network $f_{T}$ with depth $L$ and a sparse source network $f_{S}$ with depth $L\!+\!1$ be given.
@@
 3 Strong Lottery Tickets in Frozen Networks
-Although pruning at initialization provides a search space reduction, it also imposes a limitation on the sparsity range, as it can only search for SLTs with a sparsity higher than the initial pruning ratio.
-For example, an initial pruning ratio of $90\%$ vastly reduces the search space, but it limits the search to only SLTs sparser than $90\%$.
-Therefore, this reduction in search space is incompatible with exploring the whole sparsity range for optimal accuracy.
-This paper proposes a novel method that reduces the search space regardless of the desired SLT sparsity to capitalize on the training cost reduction without sacrificing accuracy.
-We explore the existence and performance of SLTs within a frozen source network, i.e., a random network that (in addition to randomly pruned parameters) has randomly chosen parameters forced to be permanent part (i.e., never pruned) of the SLT—locked parameters (see Figure 2, right).
-Since freezing is performed at initialization before the SLT search, frozen parameters do not have a score and are not eligible for exploration, thus reducing the search space.
-Locking allows to further reduce the parameter search without limiting the initial pruning ratio, reducing the search space even in the case of dense source networks.
-This section first provides a theoretical analysis of the SLT existence in such frozen networks by developing a new extension of the subset-sum approximation introduced in Section 2.
-Then, based on this analysis, it proposes partial SLT search, a novel method for finding such SLTs with a reduced search space.
-In preparation for the evaluation experiments in Section 4, we also perform a preliminary investigation of the optimal settings of the proposed method.
+The approach of finding SLTs within sparse networks by Gadhikar et al.
+(2023) offers a practical advantage: as the randomly pre-pruned parts can be reconstructed from the seed, this pre-pruning process can be exploited to reduce the memory requirement of the supermask.
+However, this approach also imposes the limitation that it can only search for SLTs with a sparsity higher than the pre-pruning ratio.
+For example, a random pre-pruning ratio of $90\%$ vastly reduces the memory size of the supermask, but it limits the search to only SLTs sparser than $90\%$.
+Therefore, their method is incompatible with exploring the whole SLT sparsity range for optimal accuracy.
+This paper proposes a novel method that allows us to find SLTs within the optimal sparsity range while still capitalizing on the compression gains offered by random connectivity initialization.
+We explore the performance of SLTs within a frozen source network, i.e., a random network that (in addition to randomly pruned parameters) has randomly chosen parameters forced to be permanent part (i.e., never pruned) of the SLT—locked parameters (see Figure 2, right).
+Since the random locking pattern can be reconstructed from seed in the same way as the random pre-pruning pattern, our freezing method allows us to compress SLTs further.
+Additionally, locking allows us to extend the benefits of pre-pruning to the scenarios where the SLT sparsity should not be too high, e.g., as found in graph neural networks (Huang et al., 2022; Yan et al., 2023).
+This section first describes the construction of frozen networks and the freezing pattern encoding for model compression.
+Then, we perform a preliminary experiment of the optimal settings for the proposed method in preparation for the evaluation experiments in Section 4.
+3.1 Partial Freezing for Enhanced SLT Compression
+Figure 3: Pre-pruning and locking set the bounds of the SLT sparsity that can be found. These optimal bounds are investigated in Section 3.2.
+In our method, since frozen regions in the source network are completely random and thus can be reconstructed from seed, we need to memorize only the optimized supermask region.
+The total amount of frozen parameters is determined by a global freezing ratio $F_{r}$, which is the sum of the respective pre-pruning ratio $P_{r}$ and locking ratio $L_{r}$.
+Therefore, we optimize the $1\!-\!F_{r}$ of the parameters (i.e., non-frozen parameters) of the original dense network.
+As explained previously, it is not possible to search for an SLT with a lower sparsity (denser) than the frozen source network.
+Consequently, as shown in Figure 3, the pre-pruning ratio $P_{r}$ sets the lower bound of the sparsity of the SLTs that can be found.
+On the other hand, it is not possible to search for an SLT that prunes more parameters than those available, so the locking ratio $L_{r}$ sets the upper bound of the sparsity of the SLTs that can be found.
+(Specifically, the upper bound is given by $1\!-\!L_{r}$.) Thus, these ratios allow us to freely control the proportion of pre-pruning and locking ratios and memory size reduction ratio for the desired SLT sparsity.
+Frozen network construction:
+We assume that the $l$-th layer of network with depth $L$ has weights $\bm{w}^{(l)}\in\mathbb{R}^{n^{(l)}}$.
+The number of frozen parameters is determined by the layer-wise freezing ratios $p_{f}^{(1)},...,p_{f}^{(L)}$, where each $p_{f}^{(l)}$ is defined as the sum of the layer-wise pre-pruning ratio $p_{p}^{(l)}$ and the layer-wise locking ratio $p_{l}^{(l)}$, and $p_{f}^{(l)},p_{p}^{(l)},p_{l}^{(l)}\in[0,1]$.
+To prune $p_{f}^{(l)}n^{(l)}$ parameters and lock $p_{l}^{(l)}n^{(l)}$ parameters, we generate two random masks, a pre-pruning mask ${\bm{m}^{(l)}_{p}\in\{0,1\}^{n^{(l)}}}$ and a locking mask ${\bm{m}^{(l)}_{l}\in\{0,1\}^{n^{(l)}}}$, so that they satisfy ${||1-\bm{m}^{(l)}_{p}||=p_{p}^{(l)}n^{(l)}}$, ${||\bm{m}^{(l)}_{l}||=p_{l}^{(l)}n^{(l)}}$, and ${(1-\bm{m}^{(l)}_{p})\cdot\bm{m}^{(l)}_{l}=0}$ (i.e., we require pre-pruning and locking to be implemented without overlap).
+The layer-wise weights frozen with these masks are calculated as ${\bm{w}^{(l)}_{f}:=(\bm{m}_{p}^{(l)}\odot(1-\bm{m}_{l}^{(l)})+\bm{m}_{l}^{(l)})\odot\bm{w}^{(l)}}$.
+These masks are fixed during training, and we search for SLTs only in the parts where ${\bm{m}_{p}^{(l)}\odot(1-\bm{m}_{l}^{(l)})}$ is one.
+Setting the layer-wise ratios:
+Our method considers the following two existing strategies for determining the layer-wise pre-pruning ratio from the desired global pre-pruning ratio of the network:
+  • Erdős-Rényi-Kernel (ERK): The pre-pruning ratio of layer $l$ is proportional to the scale ${(C_{in}^{(l)}+C_{out}^{(l)}+k_{h}^{(l)}+k_{w}^{(l)})/(C_{in}^{(l)}\cdot C_{out}^{(l)}\cdot k_{h}^{(l)}\cdot k_{w}^{(l)})}$, where $C_{in}^{(l)}$, $C_{out}^{(l)}$, $k_{h}^{(l)}$, and $k_{w}^{(l)}$ denote input channels, output channels, kernel height, and kernel width of the layer $l$, respectively (Evci et al., 2020).
+  • Edge Per Layer (EPL): Each layer’s pre-pruning ratio is set so that they all have the same number of remaining weights (Price & Tanner, 2021; Gadhikar et al., 2023).
+The same strategy is employed to determine the layer-wise pre-pruning and freezing ratios from their respective global ratios, and then the locking ratios are calculated as the difference between the corresponding freezing and pre-pruning ratios.
+Freezing pattern encoding for model compression:
+The freezing pattern can be encoded during inference as a ternary mask—a freezing mask—that indicates whether a parameter is pruned, locked, or part of the supermask.
+For example, encoding pre-pruning as $-1$, locking as $+1$, and supermask inclusion as 00, the layer-wise freezing mask can be encoded as $\bm{m}^{(l)}_{l}+(\bm{m}^{(l)}_{p}-1)\in\{-1,0,1\}^{n^{(l)}}$.
+Since this freezing mask is also random and fixed, it can be regenerated from its seed and ratios, similarly to the random weights.
+Furthermore, the supermask size is reduced by excluding from it the frozen parameters, so the SLTs found by this method can be compressed in inference to an even smaller memory size than those produced by the existing SLT literature (Hirose et al., 2022; Okoshi et al., 2022; López García-Arias et al., 2023), reducing costly off-chip memory access on specialized neural inference accelerators (Hirose et al., 2022; Chen et al., 2023), and thus offering an opportunity to perform even faster and more energy-efficient inference processing.
+3.2 Optimal Pruning:Locking Proportion for Freezing
+Figure 4: Different prune:lock proportions of a $80\%$ freezing ratio using a Conv6.
+Here, we perform a preliminary investigation of the optimal pruning:locking proportion for each given desired SLT sparsity by varying the proportion with a fixed freezing ratio of the network.
+Figure 4 explores different configurations of a $80\%$ freezing ratio—the situation where the supermask memory size is $20\%$ of the SLT in the dense source network—on a Conv6 network and compares the performance of the found SLTs on CIFAR-1010{10}10.
+As expected, the best-performing SLT of each prune:lock configuration is found at the center of the non-frozen region, where the number of candidate subnetworks is maximized.
+In other words, when the non-frozen region accounts for $S\in[0,1]$ of the entire source network, the optimal SLT sparsity is $k=P_{r}+S/2$, where $P_{r}=\sum_{l}p^{(l)}_{p}n^{(l)}/\sum_{k}n^{(k)}$ is the global pre-pruning ratio of the network.
+Conversely, for a given freezing ratio $F_{r}$ of the network and a desired SLT sparsity $k$, the optimal position of the frozen region is set by the pre-pruning ratio $P_{r}=k-(1-F_{r})/2$ and the corresponding locking ratio $L_{r}=F_{r}-P_{r}$.
+In the cases where this would result in $P_{r}\!<\!0$ or $L_{r}\!<\!0$, we choose a best-effort approach that keeps the desired freezing ratio and sets the bounds to $P_{r}\!=\!0$ or $L_{r}\!=\!0$, respectively.
+Additionally, Figure 4 compares the two strategies for setting ratios considered in Section 3.1, showing that EPL outperforms ERK in all cases.
+Consequently, hereafter, the proposed method sets the global ratios in order to position the frozen region center as close as possible to the desired SLT sparsity, and then sets the layer-wise ratios using EPL.
 3.3 SLT Existence in Frozen Networks
-Here we provide a theoretical result indicating that an SLT capable of approximating a target network exists in a frozen network.
-This result is proved by extending the subset-sum approximation lemma (Lemma 2.1) to the case where some parameters are locked (for detailed proof see Appendix A).
-Lemma 3.1 (Subset-Sum Approximation in Randomly Locked Networks (Lemma A.2)).
+One question comes to mind here: does SLT exist that approximates a given target network, even if the parameters are randomly frozen (i.e., pruned or locked) in advance?
+It has been shown by Gadhikar et al.
+(2023) that an SLT that approximates a given target network exists in the pre-pruned source network if the source network is sufficiently wider and deeper than the target, but it is not known whether such an SLT exists in the situation of a frozen source network.
+Here, we provide a rough theoretical result indicating that an SLT capable of approximating a target network exists in a frozen network with a sufficiently large width.
+This result is proved by extending the subset-sum approximation lemma (Lemma 2.1) to the case where some parameters are locked (for detailed proof, see Appendix A).
+Lemma 3.1 (Subset-Sum Approximation in Randomly Locked Networks).
 Let $X_{1},...,X_{n}$ be as in Lemma 2.1, and $M^{\prime}_{1},...,M^{\prime}_{n}\sim Ber(q)$ be independent, Bernoulli distributed random variables with $q\in(0,1)$.
 Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by the sum of $\sum_{i=1}^{n}M^{\prime}_{i}X_{i}$ and a subset-sum of $(1-M^{\prime}_{i})X_{i}$ if $n$ is sufficiently large.
-Then, by combining Lemma 3.1 and Lemma 2.2, we extend the subset-sum approximation to the situation where some random variables are frozen (i.e., pruned or locked).
-Lemma 3.2 (Subset-Sum Approximation in Frozen Networks (Lemma A.3)).
+Then, by using the same proof procedure as Lemma 3.1, we extend the subset-sum approximation to the situation where some random variables are frozen (i.e., pruned or locked).
+Lemma 3.2 (Subset-Sum Approximation in Frozen Networks).
 Let $X_{1},...,X_{n}$ be as in Lemma 2.1, $M_{1},...,M_{n}\sim Ber(p)$ be as in Lemma 2.2, and $M^{\prime}_{1},...,M^{\prime}_{n}\sim Ber(q)$ be as in Lemma 3.1.
 Then, except with exponentially small probability, any $z\in[-1,1]$ can be approximated by the sum of $\sum_{i=1}^{n}M_{i}M^{\prime}_{i}X_{i}$ and a subset-sum of $M_{i}(1-M^{\prime}_{i})X_{i}$ if $n$ is sufficiently large.
-Finally, by applying Lemma 3.2 to the Theorem 2.5 in Gadhikar et al. [10] instead of Lemma 2.2, it follows that an SLT approximating a target network exists within a frozen network.
-Theorem 3.3 (SLT Existence in Frozen Networks (Theorem A.4)).
-Let a target network $f_{T}$ with depth $L$ and a partially frozen source network $f_{S}$ with depth $L\!+\!1$ be given.
-Assume that the source network is randomly frozen with pruning ratio $p_{l}$ and locking ratio $q_{l}$ for each $l$-th layer.
-Also assume that these networks have ReLU activation function and initialized by using a uniform distribution $U[-1,1]$.
-Then, except with exponentially small probability, a subnetwork $f_{SLT}$ exists in the frozen source network $f_{S}$ such that $f_{SLT}$ approximates the target network $f_{T}$ if the width of $f_{S}$ is sufficiently large.
-Figure 4: Pruning and locking set the bounds of the search space, whose optimal position is investigated.
-3.2 Partial SLT Search in Frozen Networks
-This paper proposes a novel method for searching SLTs within frozen networks: partial SLT search.
-It only searches in part of the source network, as it has been pruned at initialization; and it only searches for part of the final SLT, as some of its parameters are predefined by locking.
-Figure 4 illustrates how network freezing reduces the search space.
-The total amount of frozen parameters is determined by a global freezing ratio $F_{r}$, which is the sum of the respective pruning ratio $P_{r}$ and locking ratio $L_{r}$.
-Therefore, the SLT search space $S$ is formed by the $1\!-\!F_{r}$ non-frozen parameters of the original dense network.
-As explained previously, it is not possible to search for an SLT with a lower sparsity (denser) than the source network.
-Consequently, the pruning ratio sets the lower bound of the search space.
-On the other hand, it is not possible to search for an SLT that prunes more parameters than those available, so the locking ratio sets the upper bound of the search space.
-(Specifically, the upper bound is given by $1\!-\!L_{r}$.) Thus, these ratios allow to freely control the size and position of the search space over the SLT sparsity spectrum.
-Frozen network construction:
-We assume that the $l$-th layer of network with depth $L$ has weights $\bm{w}^{(l)}\in\mathbb{R}^{n^{(l)}}$.
-The number of frozen parameters is determined by the layer-wise freezing ratios $p_{f}^{(1)},...,p_{f}^{(L)}$, where each $p_{f}^{(l)}$ is defined as the sum of the layer-wise pruning ratio $p_{p}^{(l)}$ and the layer-wise locking ratio $p_{l}^{(l)}$, and $p_{f}^{(l)},p_{p}^{(l)},p_{l}^{(l)}\in[0,1]$.
-To prune $p_{f}^{(l)}n^{(l)}$ parameters and lock $p_{l}^{(l)}n^{(l)}$ parameters, we generate two random masks, a pruning mask $\bm{m}^{(l)}_{p}\in\{0,1\}^{n^{(l)}}$ and a locking mask $\bm{m}^{(l)}_{l}\in\{0,1\}^{n^{(l)}}$, so that they satisfy $||1-\bm{m}^{(l)}_{p}||=p_{p}^{(l)}n^{(l)}$, $||\bm{m}^{(l)}_{l}||=p_{l}^{(l)}n^{(l)}$, and $(1-\bm{m}^{(l)}_{p})\cdot\bm{m}^{(l)}_{l}=0$ (i.e., we require pruning and locking to be implemented without overlap).
-The layer-wise weights frozen with these masks are calculated as
-$\bm{w}^{(l)}_{f}$ $:=(\bm{m}_{p}^{(l)}\odot(1-\bm{m}_{l}^{(l)})+\bm{m}_{l}^{(l)})\odot\bm{w}^{(l)}.$ (1)
-These masks are fixed during training, and we search for SLTs only in the parts where $\bm{m}_{p}^{(l)}\odot(1-\bm{m}_{l}^{(l)})$ is one.
-Setting the layer-wise ratios:
-Our method considers the following two existing strategies for determining the layer-wise pruning ratio from the desired global pruning ratio of the source network:
-  • Edge Per Layer (EPL): Each layer’s pruning ratio is set so that they all have the same number of remaining weights [29, 10].
-  • Erdős-Rényi-Kernel (ERK): The pruning ratio of layer $l$ is proportional to the scale $(C_{in}^{(l)}+C_{out}^{(l)}+k_{h}^{(l)}+k_{w}^{(l)})/(C_{in}^{(l)}\times C_{out}^{(l)}\times k_{h}^{(l)}\times k_{w}^{(l)})$, where $C_{in}^{(l)}$, $C_{out}^{(l)}$, $k_{h}^{(l)}$, and $k_{w}^{(l)}$ denote input channels, output channels, kernel height, and kernel width of the layer $l$, respectively [7].
-The same strategy is employed to determine the layer-wise pruning and freezing ratios from their respective global ratios, and then the locking ratios are calculated as the difference between the corresponding freezing and pruning ratios.
-Freezing pattern encoding for model compression:
-The freezing pattern can be encoded during inference as a ternary mask—a freezing mask—that indicates whether a parameter is pruned, locked, or part of the supermask.
-For example, encoding pruning as $-1$, locking as $+1$, and supermask inclusion as 00, the layer-wise freezing mask can be encoded as $\bm{m}^{(l)}_{l}+(\bm{m}^{(l)}_{p}-1)\in\{-1,0,1\}^{n^{(l)}}$.
-Since this freezing mask is also random and fixed, it can be regenerated from its seed and ratios, similarly to the random weights.
-Furthermore, the supermask size is reduced by excluding from it the frozen parameters, so the SLTs found by this method can be compressed in inference to an even smaller memory size than those produced by the existing SLT literature [14, 26, 21].
-Also, this memory size compression reduces costly off-chip memory access on specialized neural inference accelerators [14, 4], thus SLTs within frozen networks have the potential to perform even faster and more energy-efficient inference processing.
-3.3 Optimal Positioning of the Search Space
-Figure 5: Different prune:lock proportions of a $80\%$ freezing ratio using a Conv6.
-Here we perform a preliminary investigation of the optimal positioning of the search space bounds for each given desired SLT sparsity by exploring different pruning:locking proportions with a fixed freezing ratio of the network.
-Figure 5 explores different configurations of a $80\%$ freezing ratio—a search space size of $20\%$ of the original dense source network—on a Conv6 network and compares the performance of the found SLTs on CIFAR-1010{10}10.
-As expected, the best-performing SLT of each prune:lock configuration is found at the center of the search space, which maximizes the number of candidate subnetworks.
-In other words, within a search space size $S$, the optimal SLT sparsity is $k=P_{r}+S/2$, where $P_{r}=\frac{\sum_{l}p^{(l)}_{p}n^{(l)}}{\sum_{k}n^{(k)}}$ is the global pruning ratio of the network.
-Conversely, for a given freezing ratio $F_{r}$ of the network and a desired SLT sparsity $k$, the optimal search space position is set by the pruning ratio $P_{r}=k-(1-F_{r})/2$ and the corresponding locking ratio $L_{r}=F_{r}-P_{r}$.
-In the cases where this would result in $P_{r}\!<\!0$ or $L_{r}\!<\!0$, we choose a best-effort approach that keeps the desired freezing ratio and sets the search space bounds to $P_{r}\!=\!0$ or $L_{r}\!=\!0$, respectively.
-Additionally, Figure 5 compares the two strategies for setting ratios considered in Section 3.2, showing that EPL outperforms ERK in all cases.
-Consequently, hereafter the proposed method sets the global ratios in order to position the search space center as close as possible to the desired SLT sparsity, and then sets the layer-wise ratios using EPL.
+Finally, by applying Lemma 3.2 to the Theorem 2.5 in Gadhikar et al. (2023) instead of Lemma 2.2, it follows that an SLT approximating a target network exists within a frozen network.
+Theorem 3.1 (SLT Existence in Frozen Networks).
+Let a target network $f_{\textrm{T}}$ with depth $L$ and a partially frozen source network $f_{\textrm{S}}$ with depth $L\!+\!1$ be given.
+Assume that the source network is randomly frozen with pruning ratio $p_{l}$ and locking ratio $q_{l}$ for each $l$-th layer, except for 00-th and $1$-st layers.
+Also assume that these networks use the ReLU activation function and are initialized with a uniform distribution $U[-1,1]$.
+Then, except with exponentially small probability, a subnetwork $f_{\textrm{SLT}}$ exists in the frozen source network $f_{\textrm{S}}$ such that $f_{\textrm{SLT}}$ approximates the target network $f_{\textrm{T}}$ if the width of $f_{\textrm{S}}$ is sufficiently large.
+Note that this theorem is intended to provide an existential guarantee for the question raised at the beginning of this subsection, whether the freezing operation hurts the existence of SLTs that can approximate an arbitrary network; thus, we do not focus on the sharpness of the probability bound to keep the results simple.
+Nevertheless, proving sharper bounds may lead to capturing the effect of the freezing operation, which is left for future work.
 4 Experiments
-This section demonstrates that freezing reduces the search space in a broad range of situations by evaluating it on image classification and graph node classification.
+Figure 5: Impact of the freezing ratio on different architectures. Pruning and locking ratios are set following Section 3.2.
+This section demonstrates that freezing reduces the SLT memory size in a broad range of situations by evaluating it on image classification and graph node classification.
 We evaluate the impact of the freezing ratio on various network architectures, identifying three scenarios.
-Then, we explore trade-offs between accuracy, search space size, and model memory size for different network widths.
+Then, we explore trade-offs between accuracy and model memory size for different network widths and architectures.
 4.1 Experimental Settings
-The proposed partial SLT search is evaluated on image classification using the CIFAR-1010{10}10 [19] and ImageNet [31] datasets, and on node classification using the OGBN-Arxiv [16] dataset.
+We evaluate the SLTs within frozen networks on image classification using the CIFAR-1010{10}10 (Krizhevsky, 2009) and ImageNet (Russakovsky et al., 2015) datasets, and on node classification using the OGBN-Arxiv (Hu et al., 2020) dataset.
 CIFAR-1010{10}10 and ImageNet train data are split into training and validation sets with a 4:1 ratio, while for OGBN-Arxiv we use the default set split.
 We test the models with the best validation accuracy and report the mean of three experiment repetitions for CIFAR-1010{10}10 and OGBN-Arxiv, and the result of one experiment for ImageNet.
-Standard deviation of experiments conducted more than once is plotted as error bars.
-For image classification we employ the VGG-like Conv6 [32, 30], ResNet [13], and Wide ResNet [40] architectures, and for graph node classification the 4-layer modified GIN [37] architecture in Huang et al.
-[17], all implemented with no learned biases.
-ResNet and GIN use non-affine Batch Normalization [18], while Conv6 has no normalization layers.
-Random weights are initialized with the Kaiming Uniform distribution, while weight scores are initialized with the Kaiming Normal distribution [12].
-SLTs are searched using an extension of Edge-Popup [30] that enforces the desired sparsity globally instead of per-layer [33].
+The standard deviation of experiments conducted more than once is plotted as error bars.
+For image classification we employ the VGG-like Conv6 (Simonyan & Zisserman, 2014; Ramanujan et al., 2020), ResNet (He et al., 2016), and Wide ResNet (Zagoruyko & Komodakis, 2016) architectures, and for graph node classification the 4-layer modified GIN (Xu et al., 2019) architecture in Huang et al.
+(2022), all implemented with no learned biases.
+ResNet and GIN use non-affine Batch Normalization (Ioffe & Szegedy, 2015), while Conv6 has no normalization layers.
+Random weights are initialized with the Kaiming Uniform distribution, while weight scores are initialized with the Kaiming Normal distribution (He et al., 2015).
+SLTs are searched using an extension of Edge-Popup (Ramanujan et al., 2020) that enforces the desired SLT sparsity globally instead of per-layer (Sreenivasan et al., 2022).
 On CIFAR-1010{10}10, scores are optimized for $100$ epochs using stochastic gradient descent with momentum $0.9$, batch size $128$, weight decay $0.0001$, and initial learning rates of $0.01$ and $0.1$ for Conv6 and ResNet-1818{18}18, respectively.
 On ImageNet, scores are optimized by the same setting as ResNet-1818{18}18 on CIFAR-1010{10}10, but a $256$ batch size.
-On OBGN-Arxiv, scores are optimized for $400$ epochs using AdamW [23] with weight decay $0.0001$ and initial learning rate $0.01$.
-All experiments use cosine learning rate decay [22].
+On OBGN-Arxiv, scores are optimized for $400$ epochs using AdamW (Loshchilov & Hutter, 2019) with weight decay $0.0001$ and initial learning rate $0.01$.
+All experiments use cosine learning rate decay (Loshchilov & Hutter, 2017).
 These can be adequately verified with two NVIDIA H100 SXM5 94GB GPUs.
-Figure 6: Impact of the freezing ratio on different architectures. Pruning and locking ratios are set following Section 3.3.
-4.2 Reducing the Search Space with Freezing
-An experiment is first conducted to investigate the effect of the freezing ratio of the network on SLT accuracy. We identify three scenarios, represented in Figure 6.
+4.2 Varying Desired SLT Sparsity at Different Freezing Ratios
+This section investigates the effect of the freezing ratio of the network on desired SLT accuracy. We identify three scenarios, represented in Figure 5.
 In the cases where the optimal SLT sparsity is found at intermediate sparsity—e.g., around $50\%$ for Conv6 in LABEL:fig:sparsity_acc_conv6—pruning and locking can be applied with equally high ratios.
-Results show that the search space can be reduced by $40\%$ with small impact on accuracy, and by $70\%$ with still a moderate accuracy drop of $5$ points.
+Results show that the supermask memory size can be reduced by $40\%$ with a small impact on accuracy, and by $70\%$ with still a moderate accuracy drop of $5$ points.
 Applying the much larger ResNet-1818{18}18 to the same task results in much stronger overparametrization.
-Therefore, optimal SLTs are found in the higher sparsity range, as revealed by LABEL:fig:sparsity_acc_resnet18, benefiting of much higher pruning than locking.
-Even though $90\%$ of the search space is reduced in this scenario, we can find $90\%$ sparse SLTs with $88.1\%$ accuracy.
+Therefore, optimal SLTs are found in the higher sparsity range, as revealed by LABEL:fig:sparsity_acc_resnet18, benefiting from much higher pruning than locking.
+Even though $90\%$ of the memory size is reduced in this scenario, we can find $90\%$ sparse SLTs with $88.1\%$ accuracy.
 As an example of the scenario where optimal SLT sparsities are found in the denser range, benefiting from a higher locking ratio, we evaluate GIN in LABEL:fig:sparsity_acc_gin.
-Compared with the best-performing SLT found in the dense GIN, of $70.1\%$ accuracy and $20\%$ sparsity, by freezing $50\%$ of the search space our method finds similarly performing SLTs of $69.8\%$ accuracy with $40\%$ sparsity.
-Interestingly, with low SLT sparsity (e.g., $10\%$ sparsity) in all three scenarios, despite the reduced search space, SLTs within frozen networks achieve higher accuracy than SLTs within dense networks.
-These results suggest that parameter freezing at an appropriate ratio has the effect of avoiding the inclusion of low-grade local optimal solutions in the search space.
-It is known that Edge-Popup struggles to find the SLTs of the global optimal solution [9].
-While searching for SLTs in a dense network leads to convergence to a local optimal solution, a moderate reduction of the search space may reduce the number of less accurate local optimal solutions and facilitate convergence to a more accurate local optimal solution within the reduced search space.
+Compared with the best-performing SLT found in the dense GIN, of $70.1\%$ accuracy and $20\%$ sparsity, by freezing $50\%$ of the memory size, our method finds similarly performing SLTs of $69.8\%$ accuracy with $40\%$ sparsity.
+Interestingly, with low SLT sparsity (e.g., $10\%$ sparsity) in all three scenarios, despite the reduced parameters to be optimized, SLTs within frozen networks achieve higher accuracy than SLTs within dense networks.
+These results imply that parameter freezing at an appropriate ratio has the effect of avoiding the inclusion of low-grade local optimal solutions in the search space.
+While searching for SLTs in a dense network by Edge-Popup leads to convergence to a local optimal solution (Fischer & Burkholz, 2022), a moderate random parameter freezing may reduce the number of less accurate local optimal solutions and facilitate convergence to a more accurate local optimal solution within the reduced search space.
 In other words, we conjecture that if the network is properly frozen, the local optimal solution for the reduced search space is close to the global optimal solution of the entire search space.
-4.3 Accuracy-to-Search Space Trade-Off
-Section 3 sets as a requisite of the subset-sum proof that the source network has a sufficiently larger width than the target network.
-Figure 6 explores varying the width of the source network to analyze its impact on accuracy and search space size.
+Figure 6: Compared to sparse ($\bullet$) or dense ($\bullet$) source networks, freezing achieves better accuracy-to-model memory size trade-off (top-left is better).
+4.3 Accuracy-to-Model Memory Size Trade-Off
+The freezing mask compression scheme proposed in Section 3.1 allows to reduce the model memory size during inference by regenerating both the random weights and the freezing mask with random number generators.
+Here we consider this compression and investigate the accuracy-to-model size trade-off offered by the SLTs found in the frozen networks.
+Model size refers to the total memory size of model parameters that need to be stored.
+The random weights and the frozen parts of the supermask are excluded, since they can be regenerated from seed, whereas each non-frozen element of the supermask and each learned batchnorm parameter are counted as 1 and 32 bits, respectively.
+Furthermore, we also compare the Kaiming Uniform (KU) weight initialization used so far with the binary weights provided by the Signed Kaiming Constant (SKC) initialization (Ramanujan et al., 2020), which can be exploited for reduced computational cost in neural engines (Hirose et al., 2022).
+For SKC, we scale weights by $1/\sqrt{1-k_{l}}$, where $k_{l}$ is the sparsity of each layer, as proposed by Ramanujan et al.
+(2020).
+Note that since we are interested in the accuracy of the SLTs compressed as much as possible obtained by each method, the pre-pruning ratio for the pruning-only case is determined as $(\textrm{target sparsity}-5)\%$ to compress the SLT as much as possible while allowing to search for SLTs.
+On the other hand, we evaluate freezing with a higher compression ratio than pruning only.
+This setting means we evaluate our method under more strict conditions than pruning-only from the compression ratio perspective.
+Figure 6 explores varying the width of the source network to analyze its impact on accuracy and model size.
 SLT sparsity is fixed to that of the best performing SLT found in a dense source network in Figure 5: $50\%$ in Conv6, $90\%$ in ResNet-1818{18}18, and $20\%$ in GIN.
-The results show in all cases that the found SLTs in wide frozen networks for a given sparsity achieve similar or higher performance with a similar or smaller search space size compared to the dense source network.
-That is, the proposed method improves the accuracy-to-search space size trade-off.
-As shown in LABEL:fig:search_space_acc_conv6 and LABEL:fig:search_space_acc_gnn, in the scenarios that only allow for medium to low pruning, there is a limit to the number of parameters that can be pruned for a given SLT sparsity, and only aggressive pruning leads to less accurate SLTs.
-By additionally locking parameters, our method reduces the search space further and finds SLTs of higher accuracy than those within sparse or dense networks for a similar search space size.
-Pruning only improves the trade-off in the scenario that allows higher pruning than locking, as shown in LABEL:fig:search_space_acc_resnet18, while our method improves the accuracy-to-search space trade-off in all scenarios.
-Figure 7: Freezing achieves better accuracy-to-search space trade-off compared to sparse ($\bullet$) or dense ($\bullet$) source networks.
-SLT sparsity set according to highest accuracy in Figure 6 with dense source networks.
-Figure 8: Compared to sparse ($\bullet$) or dense ($\bullet$) source networks (top-left is better), freezing achieves better accuracy-to-memory size trade-off.
-4.4 Accuracy to Model Memory Size Trade-Off
-The freezing mask compression scheme proposed in Section 3.2 allows to reduce the memory model size during inference by regenerating both the random weights and the freezing mask with a random number generator (RNG).
-Here we consider this compression and investigate the accuracy-to-model size trade-off offered by the SLTs found by the proposed method.
-Furthermore, we also compare the Kaiming Uniform (KU) weight initialization used so far with the binary weights provided by the Signed Kaiming Constant (SKC) initialization [30], which can be exploited for reduced computational cost in neural engines [14].
-For SKC we scale weights by $1/\sqrt{1-k_{l}}$, where $k_{l}$ is the sparsity of each layer, as proposed by Ramanujan et al.
-[30].
-As illustrated in Figure 8, compared to the SLTs found in dense or sparse source networks, SLTs found by our method achieve similar or higher accuracy for similar or smaller model size, thus improving the accuracy-to-model size trade-off in all scenarios.
-In other words, partial SLT search not only reduces the search space during training, but also reduces the model size during inference, further improving performance over conventional SLT search methods.
-Empirically, it is known that SLTs within a SKC initialized dense network achieve better performance than with continuous random weights [30, 26, 21, 38].
+Compared to the SLTs found in dense or sparse source networks, SLTs found by our method achieve similar or higher accuracy for similar or smaller model size, thus improving the accuracy-to-model size trade-off in all scenarios.
+Empirically, it is known that SLTs within an SKC-initialized dense network achieve better performance than with continuous random weights (Ramanujan et al., 2020; Okoshi et al., 2022; López García-Arias et al., 2023; Yan et al., 2023).
 Our results show that such a trend can also be observed in frozen source networks.
 Nonetheless, in all source networks, we find that this improvement is smaller the wider the source network is, suggesting that the requirement for source networks of larger width is weaker in the case of binary weights.
-4.5 ImageNet Experiments
-Figure 9: Freezing enhances the accuracy-to-model size trade-off even on a practical setting with large-scale dataset and models.
-Finally, we evaluate our method using larger models on a large-scale dataset: deeper and wider ResNets on ImageNet.
-Since SLTs with $80\%$ sparsity achieve the highest accuracy in a dense source ResNet-5050{50}50 (for details, see Figure 10 in Appendix B), we compare the three methods using $80\%$ SLT sparsity.
-Figures 2 and 9 compare the accuracy-to-search space and accuracy-to-model size trade-offs, respectively, in finding the $80\%$ sparsity SLTs between the proposed and conventional methods with SKC using ImageNet.
-In both of the figures, such SLTs searched from non-frozen (blue), $70\%$ pruned (orange), and $65\%$ pruned and $5\%$ locked ($70\%$ frozen, green) are compared.
-Despite the more challenging setting and the significant $70\%$ search space reduction, our method finds SLTs that are substantially more accurate than other methods for the same search space size (Figure 2) and for the same model memory size (Figure 9).
-This result demonstrates that the effective combination of parameter pruning and locking at initialization can improve the SLT search efficiency even on large-scale datasets and models.
-Notably, Figure 9 reveals that the proposed method enhances the accuracy-to-model size trade-off, validating our claim it could be deployed on specialized inference hardware, such as the one like Hirose et al.
-[14], to enhance the efficiency of practical SLT applications.
-5 Limitations
-Our results have the following limitations: 1) Although the proposed method can effectively reduce the search space for training and the memory model size for inference in principle, actual hardware implementation remains for future work.
-The previous work of the SLT-specialized neural engine, Hiddenite [14], would be a promising direction for such future implementation.
-2) This paper provides theoretical support for the existence of SLTs in frozen networks, but it does not provide their superiority in the approximation capability compared to random pruning, a special case of random freezing, which may depend on the sparsity nature of target networks.
-Also, the theoretical consideration of the appropriate proportion of these ratios is left for future work.
-3) Even though we demonstrate our method on various model architectures including CNNs, ResNet families and GIN, following previous SLT work [41, 30, 21, 17, 38], we have to leave it for future work to apply our method to Transformers [35] because SLT itself has not yet been established for Transformers.
-6 Conclusion
-This paper capitalizes on the fact that pruning a randomly weighted network at initialization reduces the SLT search space, but identifies that doing so limits the search to a suboptimal sparsity region.
-This problem is tackled by freezing (i.e., pruning or locking) some parameters at initialization, excluding them from the search.
-Freezing allows to search for SLTs efficiently in the optimal sparsity range, while further reducing the search space during training and the model size during inference.
-Experimental results show that SLTs found in frozen networks improve the accuracy-to-search space and accuracy-to-model size trade-offs compared to SLTs found in dense [30, 33] or sparse source networks [10].
-Interestingly, although for weight training random parameter locking has only been found useful for reducing accuracy degradation in network compression [41, 36], we identify scenarios in SLT training where they can be used for raising accuracy.
-Our method can be interpreted as being capable of generating more useful SLT information from a random seed than previous methods, offering an opportunity for reducing off-chip memory access in specialized SLT accelerators [14].
-We hope this work will serve as a base for raising the efficiency of SLT training and inference.
-Acknowledgments and Disclosure of Funding
-This work was supported in part by JSPS KAKENHI Grant Number JP23H05489.
-Appendix A Proof for Strong Lottery Tickets (SLTs) Existence in Frozen Networks
-This section describes the proofs of the lemma and theorem introduced in the manuscript. For reference, we present a theorem presented by Gadhikar et al. [10] in advance.
-Theorem A.1 (SLT Existence in Sparse Networks [10]).
-Let input data $\mathcal{D}$ , a target network $f_{T}$ with depth $L$, and a source network $f_{S}$ with depth $L+1$, parameters $\bm{\theta}_{S}$, and edge probabilities $p_{l}$ be given.
-Assume that these networks have ReLU activation function, and each element of $\bm{\theta}_{S}$ is uniformly distributed over $[-1,1]$.
-Then, with probability at least $1-\delta$, there exists a binary mask $\bm{m}_{S}$ so that each output component $i$ is approximated as $\max_{\bm{x}\in\mathcal{D}}\left\|f_{T,i}(\bm{x})-f_{S,i}(\bm{x};\bm{\theta}_{S}\odot\bm{m}_{S})\right\|<\epsilon$ if
-$n_{S,l}\geq C\frac{n_{T,l}}{\log(1/1-p_{l+1})}\log\left(\frac{1}{\min\{\epsilon_{l},\delta/\rho\}}\right)$ (2)
-for $l\geq 1$,where
-$\epsilon_{l}$ $\coloneqq\frac{\epsilon}{n_{T,L}L}\left[(1+B_{l-1}\left(1+\frac{\epsilon}{L}\right)\prod_{k=l+1}^{L-1}\left(\left\|\bm{\theta}_{T}^{(l)}\right\|_{\infty}+\frac{\epsilon}{L}\right)\right]^{-1},$ (3)
-$B_{l}$ $\coloneqq\sup_{\bm{x}\in\mathcal{D}}\left\|\bm{x}_{T}^{(l)}\right\|_{1},$ (4)
-$\rho$ $=\frac{CN_{T}^{1+\gamma}}{\log\left(1/(1-\min_{l}\{p_{l}\})\right)^{1+\gamma}}\log\left(\frac{1}{\min\{\min_{l}\{\epsilon_{l}\},\delta\}}\right)$ (5)
-for any $\gamma\geq 0$. Moreover, we need to satisfy
-$n_{S,0}\geq Cd\log\left(\frac{1}{\min\left\{\epsilon_{1},\delta/\rho\right\}}\right),$ (6)
-where $C>0$ denotes a generic constant that is independent of $n_{T,l},L,p_{l},\delta$, and $\epsilon$.
-We first extend the lemma of the subset-sum approximation shown by Lueker [24] to the case where some random variables are always included in the subset-sum.
-Lemma A.2 (Subset-Sum Approximation in Randomly Locked Networks).
-Let $X_{1},\cdots,X_{n}\sim U(-1,1)$ be independent, uniformly distributed random variables, and $M_{1},\cdots,M_{n}\sim\mathrm{Ber}(q)$ be independent, Bernoulli distributed random variables with $q\in(0,1)$.
-Let $\varepsilon,\delta>0$.
-Then, with probability at least $1-\delta$, for any $z\in[-1,1]$, there exists indices $I\subset\{1,\cdots,n\}$ such that $|z-\sum_{i=1}^{n}M_{i}X_{i}-\sum_{i\in I}(1-M_{i})X_{i}|\leq\varepsilon$ if
-$n\geq C\log\left(\frac{8}{\varepsilon\delta}\right),$ (7)
-where $C>0$ is a constant.
-Proof.
-Let $M_{1},\cdots,M_{n}\sim\mathrm{Ber}(q)$ and $m:=\sum_{i}M_{i}$. By Hoeffding’s inequality, we have
-$\mathbb{P}\left(|m-qn|\leq\varepsilon_{M}n\right)\geq 1-2\exp\left(-\frac{\varepsilon_{M}^{2}n}{2q(1-q)}\right)$ (8)
-for $\varepsilon_{M}>0$. Thus, if we set $n\geq\frac{2q(1-q)\log(2/\delta)}{\varepsilon_{M}^{2}}$ and $\varepsilon_{M}:=\frac{\min(q,1-q)}{2}$, we have
-$(2q-\beta)n\leq m\leq\beta n$ (9)
-with $\beta:=q+\varepsilon_{M}=\min(\frac{3q}{2},\frac{1+q}{2})\in(0,1)$, with probability at least $1-\delta$.
-In particular, the number of non-vanishing terms in the sum $\sum_{i\in I}(1-M_{i})X_{i}$ is $n-m\geq(1-\beta)n$ as long as each $X_{i}$ is non-zero.
-Now fix $M_{1},\cdots,M_{n}\sim\mathrm{Ber}(q)$ with $(2q-\beta)n\leq m\leq\beta n$.
-The goal is to approximate $z\in[-1,1]$ and $\sum_{i=1}^{n}M_{i}X_{i}$ by subset sum from $\{(1-M_{i})X_{i}:M_{i}=0\}$.
-For simplicity, we split it into two parts:
-$\{(1-M_{i})X_{i}:M_{i}=0\}=\{X^{\prime}_{1},\cdots,X^{\prime}_{n_{1}}\}\cup\{X^{\prime\prime}_{1},\cdots,X^{\prime\prime}_{n_{2}}\},$ (10)
-where the former part is used for approximating $z\in[-1,1]$ and the latter part for approximating $\sum_{i=1}^{n}M_{i}X_{i}$.
-To approximate $z\in[-1,1]$, we can directly apply Corollary 2.5 from Lueker [24]:
-$\mathbb{P}\left(\forall z\in\left[-1,1\right],\exists I\subset\{1,\cdots,n_{1}\}\text{ s.t. }|z-\sum_{i\in I}X^{\prime}_{i}|\geq\varepsilon\right)\geq 1-\delta$ (11)
-whenever $n_{1}\geq C\log(2/\varepsilon\delta)$.
-To approximate $\sum_{i=1}^{n}M_{i}X_{i}$, we have to evaluate its norm. By Hoeffding’s inequality on $X_{i}$’s with $M_{i}\not=0$, we have
-$\mathbb{P}\Big{(}\Big{|}\sum_{i=1}^{n}M_{i}X_{i}\Big{|}\leq\alpha m\Big{)}\geq 1-2\exp\left(-\frac{3\alpha^{2}m}{2}\right),$ (12)
-for any fixed $\alpha>0$, whose value will be specified later.
-Thus $\Big{|}\sum_{i=1}^{n}M_{i}X_{i}\Big{|}\leq\alpha m\leq\alpha\beta n$ holds with probability at least $1-\delta$ whenever $m\geq\frac{2\log(2/\delta)}{3\alpha^{2}}$.
-Since $m\leq\beta n$ holds, $n\geq\frac{2\log(2/\delta)}{3\alpha^{2}\beta}$ is enough.
-From the proof of Corollary 3.1 in Lueker [24], for any $\gamma\in(0,\frac{1}{4})$, we know that
-$\mathbb{P}\left(\forall z\in\left[-\gamma n_{2},\gamma n_{2}\right],\exists I\subset\{1,\cdots,n_{2}\}\text{ s.t. }|z-\sum_{i\in I}X^{\prime\prime}_{i}|\geq\varepsilon\right)$ (13)
-$\geq 1-\delta-2\exp\left(-\frac{(1-4\gamma)^{2}n_{2}}{64}\right)$ (14)
-whenever $n_{2}\geq C\log(\frac{2}{\varepsilon\delta})$. Thus if
-$n_{2}\geq\max\left(C\log(\frac{2}{\varepsilon\delta}),\frac{C^{\prime}}{(1-4\gamma)^{2}}\log(\frac{2}{\delta})\right)$ (15)
-holds, we can approximate any $z\in[-\gamma n_{2},\gamma n_{2}]$ by subset sum of $X^{\prime\prime}_{1},\cdots,X^{\prime\prime}_{n_{2}}$ with probability $1-2\delta$.
-Now we assume $(\alpha\beta/\gamma)n\leq n_{2}$.
-For example, if we set $\alpha=\frac{1}{10}$ and $\gamma=\frac{1}{5}$, the assumption is satisfied when we split $\{(1-M_{i})X_{i}\}$ by $n_{1}=n_{2}=\frac{n}{2}$ in (10).
-Then the approximation (13) can be applied to $z=\sum_{i=1}^{n}M_{i}X_{i}$ with high probability since $|\sum_{i=1}^{n}M_{i}X_{i}|\leq\alpha m\leq\alpha\beta n\leq\gamma n_{2}$ holds.
-By combining (11), (12), (15) with the fact that $n=m+n_{1}+n_{2}$, and replacing $\delta$ with $\delta/4$, we obtain the desired results. ∎
-By similar arguments as in Lueker [24] or Pensia et al.
-[28], we can easily generalize Lemma A.2 to the case where the distribution followed by $X_{i}$ contains the uniform distribution.
-Also, by combining our proof and the proof of Lemma 2.4 in Gadhikar et al.
-[10], we can prove the following extension of Lemma A.2:
-Lemma A.3 (Subset-Sum Approximation in Frozen Networks.).
-Let $X_{1},...,X_{n}$ be independent, uniformly distributed random variables so that $X\sim U(-1,1)$, $M_{1},...,M_{n}$ be independent, Bernoulli distributed random variables so that $M_{i}\sim Ber(p)$ for $p\in(0,1)$, and $M^{\prime}_{1},...,M^{\prime}_{n}$ be independent, Bernoulli distributed random variables so that $M^{\prime}_{i}\sim Ber(q)$ for $q\in(0,1)$.
-Let $\epsilon,\delta\in(0,1)$ given.
-Then for any $z\in[-1,1]$ there exists a subset $I\subseteq\{1,...,n\}$ so that with probability at least $1-\delta$ we have $\left|z-\sum_{i=1}^{n}M_{i}M^{\prime}_{i}X_{i}-\sum_{i\in I}M_{i}\left(1-M^{\prime}_{i}\right)X_{i}\right|\leq\epsilon$ if
-$n\geq C\log\left(\frac{C^{\prime}}{\varepsilon\delta}\right),$ (16)
-where $C,C^{\prime}>0$ is a constant.
-Theorem A.4 (SLT Existence in Frozen Networks).
-Let $\mathcal{D}$ be input data, $f_{T}$ be a target network with depth $L$, and $f_{S}$ be a source network with depth $L+1$, $\bm{\theta}_{S}$ be a parameter of $f_{S}$, and $p_{l},q_{l}\in(0,1)$.
-Assume that these networks have ReLU activation function, and each element of $\bm{\theta}_{S}$ is uniformly distributed over $[-1,1]$.
-Also assume that $\bm{\theta}_{S}$ is randomly pruned and locked with pruning ratio $p_{l}$ and locking ratio $q_{l}$ for each $l$-th layer.
-Then, with probability at least $1-\delta$, there exists a binary mask $\bm{m}_{S}$ so that each output component $i$ is approximated as $\max_{\bm{x}\in\mathcal{D}}\left\|f_{T,i}(\bm{x})-f_{S,i}(\bm{x};\bm{\theta}_{S}\odot\bm{m}_{S})\right\|<\epsilon$ if
-$n_{S,l}$ $\geq C_{1}\log\left(\frac{C_{2}}{\min\{\epsilon_{l},\delta/\rho\}}\right)n_{T,l},$ $(l\geq 1)$ (17)
-$n_{S,0}$ $\geq C_{3}\log\left(\frac{C_{4}}{\min\left\{\epsilon_{1},\delta/\rho\right\}}\right)d$ $(l=0)$ (18)
-where $C_{i}>0$ are constants that include $p_{l},q_{l}$, and $\rho,\epsilon_{l}$ are as defined in Theorem A.1.
-Proof.
-By replacing Lemma 2.4 in Gadhikar et al. [10] with our Lemma A.3, we obtain the desired result as an extension of Theorem 2.5 in Gadhikar et al. [10]. ∎
-Appendix B Additional Experimental Results
-B.1 Accuracy of SLT within Dense ResNet-50
-This section introduces the preliminary experiment used for determining the main experimental setup.
-Figure 10 compares the accuracy of SLTs within a dense ResNet-5050{50}50 source at different sparsity using ImageNet.
-The SLT with $80\%$ sparsity is the most accurate, reaching $66.8\%$ accuracy.
-Figure 10: Accuracy comparison of SLTs of different sparsity within a dense ResNet-5050{50}50 on ImageNet.
-B.2 Result Analysis
-Table 1 summarizes the presented results and compares the proposed method to weight training and SLT search in dense and sparse source networks for the same network architecture and SLT sparsity.
-In the case of Conv6, compared to weight training our method provides reductions of 2×2$\times$2 × and 66.4×66.4$\times$66.4 × of the search space and model size, respectively, in exchange of a small accuracy drop.
-Compared to the SLT found in the sparse source network, the SLT found by partial search achieves much higher accuracy.
-[…]
-Interestingly, the SLT found in the frozen GIN model achieves comparable accuracy to the trained-weight network, even though the model size and search space are reduced by 40.3×40.3$\times$40.3 × and 1.7×1.7$\times$1.7 ×, respectively.
-The SLT found within a dense network also achieves comparable accuracy to the trained-weight network, but the SLT found by our method is smaller.
-Generally, graph neural networks (GNNs) suffer from a generalization performance degradation due to the over-smoothing problem [20].
-Searching for SLTs within dense GNNs has been shown to mitigate the over-smoothing problem [17] and achieve higher accuracy.
-As our method is similarly accurate compared to dense networks, it is possible that parameter freezing offers an even stronger solution to the over-smoothing problem.
-Although the ImageNet experiment falls into the scenario where a relatively sparse SLT is more accurate (like the ResNet-1818{18}18 experiment on CIFAR-1010{10}10), the SLT found in a sparse network is significantly less accurate.
-On the other hand, the inclusion of parameter locking shows a significant improvement in SLT accuracy.
-These results suggest that the number of highly accurate SLT patterns decreases as the difficulty of the problem increases, and the way the strategy of search space reduction affects the performance of the found SLTs more severely.
-As also shown in Figure 2 in the main manuscript, the SLT found in the frozen Wide ResNet-5050{50}50 achieves $1.4$ points higher accuracy than the SLT found in the dense ResNet-3434{34}34, even though the search space is smaller, indicating superiority of our method in improving the accuracy-to-search space trade-off.
 Table 1: Comparison between trained-weight networks and SLTs found in dense, sparse, and frozen networks.
@@
 Acc. [%]
-Search Space
-[# Param.]
 Model
 Size [MB]
-Weight Training KU - - - $87.1$ $2.26$ M $8.63$
-SLT (Dense) SKC $50$ 00 00 $86.2$ $2.26$ M $0.27$
-SLT (Sparse) SKC $50$ $45$ 00 $66.7$ $1.24$ M $0.15$
-SLT (Frozen) SKC $50$ $25$ $25$ $84.8$ $1.13$ M $0.13$
+Weight Training KU - - - $87.1$ $8.63$
+SLT (Dense) SKC $50$ 00 00 $86.2$ $0.27$
+SLT (Sparse) SKC $50$ $45$ 00 $66.7$ $0.15$
+SLT (Frozen) SKC $50$ $25$ $25$ $84.8$ $0.13$
 ResNet-${18}$ CIFAR-1010{10}10
-Weight Training KU - - - $92.4$ $11.16$ M $42.63$
-SLT (Dense) SKC $90$ 00 00 $91.0$ $11.16$ M $1.37$
-SLT (Sparse) SKC $90$ $85$ 00 $89.9$ $1.67$ M $0.24$
-SLT (Frozen) SKC $90$ $82.5$ $2.5$ $89.8$ $1.67$ M $0.24$
-SLT (Frozen) SKC $90$ $85$ $5$ $88.9$ $1.12$ M $0.17$
+Weight Training KU - - - $92.4$ $42.63$
+SLT (Dense) SKC $90$ 00 00 $91.0$ $1.37$
+SLT (Sparse) SKC $90$ $85$ 00 $89.9$ $0.24$
+SLT (Frozen) SKC $90$ $82.5$ $2.5$ $89.8$ $0.24$
+SLT (Frozen) SKC $90$ $85$ $5$ $88.9$ $0.17$
 GIN & OGBN-Arxiv
-Weight Training KU - - - $70.1$ $0.38$ M $1.452$
-SLT (Dense) SKC $20$ 00 00 $70.0$ $0.38$ M $0.054$
-SLT (Sparse) SKC $20$ $15$ 00 $62.3$ $0.32$ M $0.047$
-SLT (Frozen) SKC $20$ 00 $40$ $69.2$ $0.23$ M $0.036$
+Weight Training KU - - - $70.1$ $1.452$
+SLT (Dense) SKC $20$ 00 00 $70.0$ $0.054$
+SLT (Sparse) SKC $20$ $15$ 00 $62.3$ $0.047$
+SLT (Frozen) SKC $20$ 00 $40$ $69.2$ $0.036$
 ResNet-${50}$ ImageNet
-Weight Training KU - - - $74.4$ $25.50$ M $97.49$
-SLT (Dense) SKC $80$ 00 00 $66.8$ $25.50$ M $3.24$
-SLT (Sparse) SKC $80$ $70$ 00 $47.1$ $7.65$ M $1.11$
-SLT (Frozen) SKC $80$ $65$ $5$ $55.2$ $7.65$ M $1.11$
+Weight Training KU - - - $74.4$ $97.49$
+SLT (Dense) SKC $80$ 00 00 $66.8$ $3.24$
+SLT (Sparse) SKC $80$ $70$ 00 $47.1$ $1.11$
+SLT (Frozen) SKC $80$ $65$ $5$ $55.2$ $1.11$
 ResNet-${34}$ ImageNet
-SLT (Dense) SKC $80$ 00 00 $62.5$ $21.78$ M $2.66$
-SLT (Sparse) SKC $80$ $70$ 00 $40.7$ $6.53$ M $0.84$
-SLT (Frozen) SKC $80$ $65$ $5$ $51.2$ $6.53$ M $0.84$
+SLT (Dense) SKC $80$ 00 00 $62.5$ $2.66$
+SLT (Sparse) SKC $80$ $70$ 00 $40.7$ $0.84$
+SLT (Frozen) SKC $80$ $65$ $5$ $51.2$ $0.84$
 ResNet-${18}$ ImageNet
-SLT (Dense) SKC $80$ 00 00 $54.0$ $11.68$ M $1.43$
-SLT (Sparse) SKC $80$ $70$ 00 $25.2$ $3.50$ M $0.45$
-SLT (Frozen) SKC $80$ $65$ $5$ $39.4$ $3.50$ M $0.45$
+SLT (Dense) SKC $80$ 00 00 $54.0$ $1.43$
+SLT (Sparse) SKC $80$ $70$ 00 $25.2$ $0.45$
+SLT (Frozen) SKC $80$ $65$ $5$ $39.4$ $0.45$
 Wide ResNet-${50}$ ImageNet
-SLT (Dense) SKC $80$ 00 00 $70.8$ $68.81$ M $8.46$
-SLT (Sparse) SKC $80$ $70$ 00 $57.7$ $20.64$ M $2.72$
-SLT (Frozen) SKC $80$ $65$ $5$ $63.9$ $20.64$ M $2.72$
-Appendix C Impact Statements
+SLT (Dense) SKC $80$ 00 00 $70.8$ $8.46$
+SLT (Sparse) SKC $80$ $70$ 00 $57.7$ $2.72$
+SLT (Frozen) SKC $80$ $65$ $5$ $63.9$ $2.72$
+4.4 ImageNet Experiments
+Finally, we evaluate our method using larger models on a large-scale dataset: deeper and wider ResNets on ImageNet.
+Since SLTs with $80\%$ sparsity achieve the highest accuracy in a dense source ResNet-5050{50}50 (for details, see Section B.1), we compare the three methods using $80\%$ SLT sparsity.
+Figure 2 compares the accuracy-to-model size trade-off in finding the $80\%$ sparsity SLTs between the proposed and conventional methods with SKC using ImageNet.
+Despite the more challenging setting and the significant $70\%$ memory size reduction, our method (green) finds SLTs that are more accurate than pre-pruning-only methods (orange) for the same model size.
+This result demonstrates that the effective combination of parameter pruning and locking at initialization can improve the SLT memory efficiency even on large-scale datasets and models.
+4.5 Result Analysis
+Table 1 summarizes the presented results and compares the proposed method to weight training and SLT search in dense and sparse source networks for the same network architecture and SLT sparsity.
+As described in Section 4.3, the model size is calculated as the sum of the unfrozen partial supermask (1 bit/parameter) and the batch normalization parameters (32 bits/parameter).
+Also, as described in Section 4.2, the target SLT sparsity is fixed to that of the best-performing SLT found in a dense source network, and we determine the pre-pruning ratio in sparse network as $(\textrm{SLT sparsity}-5)\%$ for high compression.
+In the case of Conv6, compared to weight training our method provides reductions of 66.4×66.4$\times$66.4 × of the model size, in exchange of a small accuracy drop.
+Compared to the SLT found in the sparse source network, the SLT found in the frozen source network achieves much higher accuracy.
+[…]
+Interestingly, the SLT found in the frozen GIN model achieves comparable accuracy to the trained-weight network, even though the model size is reduced by 40.3×40.3$\times$40.3 ×.
+The SLT found within a dense network also achieves accuracy comparable to that of the trained-weight network, but the SLT found by our method is smaller.
+Generally, graph neural networks (GNNs) suffer from a generalization performance degradation due to the over-smoothing problem Li et al.
+(2018).
+Searching for SLTs within dense GNNs has been shown to mitigate the over-smoothing problem Huang et al.
+(2022) and achieve higher accuracy.
+As our method is similarly accurate compared to dense networks, it is possible that parameter freezing offers an even stronger solution to the over-smoothing problem.
+Although the ImageNet experiment falls into the scenario where a relatively sparse SLT is more accurate (like the ResNet-1818{18}18 experiment on CIFAR-1010{10}10), the SLT found in a sparse network is significantly less accurate.
+On the other hand, the inclusion of parameter locking shows a significant improvement in SLT accuracy.
+These results suggest that the number of highly accurate SLT patterns decreases as the difficulty of the problem increases, and the pruning:locking proportion affects the performance of the found SLTs more severely.
+5 Limitations
+Our results have the following limitations: 1) Although the proposed method can effectively reduce the model memory size for inference in principle, actual hardware implementation remains for future work.
+The previous work of the SLT-specialized neural engine, Hiddenite Hirose et al.
+(2022), would be a promising direction for such future implementation.
+2) This paper provides theoretical support for the existence of SLTs in frozen networks, but it does not provide their superiority in the approximation capability compared to random pruning, a special case of random freezing, which may depend on the sparsity nature of target networks.
+Also, the theoretical consideration of the appropriate proportion of these ratios is left for future work.
+3) Even though we demonstrate our method on various model architectures including CNNs, ResNet families and GIN in Section 4, following previous SLT work (Zhou et al., 2019; Ramanujan et al., 2020; López García-Arias et al., 2023; Huang et al., 2022; Yan et al., 2023), we have to leave it for future work to apply our method to Transformers (Vaswani et al., 2017) because SLT itself has not yet been established for Transformers.
+6 Conclusion
+This paper capitalizes on the fact that pre-pruning a randomly weighted network reduces the supermask memory size, but identifies that doing so limits the search to a suboptimal sparsity region.
+This problem is tackled by freezing (i.e., pruning or locking) some parameters at initialization, excluding them from the search.
+Freezing allows to the search for SLTs in the optimal sparsity region, while further reducing the model size.
+Experimental results show that SLTs found in frozen networks improve the accuracy-to-model size trade-off compared to SLTs found in dense (Ramanujan et al., 2020; Sreenivasan et al., 2022) or sparse networks (Gadhikar et al., 2023).
+Interestingly, although for weight training random parameter locking has only been found useful for reducing accuracy degradation in network compression (Zhou et al., 2019; Wimmer et al., 2020), we identify scenarios in SLT training where they can be used for raising accuracy.
+Our method can be interpreted as being capable of generating more useful SLT information from a random seed than previous methods, offering an opportunity for reducing off-chip memory access in specialized SLT accelerators (Hirose et al., 2022).
+Additionally, the reduced number of parameters to be optimized may be exploited for training cost reduction.
+Appropriate freezing may improve the SLT search efficiency, including preventing convergence to local low-precision solutions as experimentally confirmed in Section 4.2, and improving convergence speed through reduced parameters to be optimized.
+Investigations about the SLT search efficiency remain as future work.
+Acknowledgments
+This work was supported in part by JSPS KAKENHI Grant Number JP23H05489 and JST-ALCA-Next Japan Grant # JPMJAN24F3.
+Impact Statements
 This paper presents work whose goal is to advance the field of machine learning.
-Among the many potential societal consequences of our work, we highlight its potential to reduce the computational cost of inference and training of neural networks, and thus their energy consumption footprint.
+Among the many potential societal consequences of our work, we highlight its potential to reduce the computational cost of inference of neural networks, and thus their energy consumption footprint.
+Appendix A Proof for Strong Lottery Tickets (SLTs) Existence in Frozen Networks
+This section describes the proofs of the lemma and theorem introduced in the manuscript.
+We first extend the lemma of the subset-sum approximation shown by Lueker (1998) to the case where some random variables are always included in the subset-sum.
+Lemma A.1 (Subset-Sum Approximation in Randomly Locked Networks).
+Let $X_{1},\cdots,X_{n}\sim U(-1,1)$ be independent, uniformly distributed random variables, and $M_{1},\cdots,M_{n}\sim\mathrm{Ber}(q)$ be independent, Bernoulli distributed random variables with $q\in(0,1)$.
+Let $10<\varepsilon,\delta<$ 1.
+Then, with probability at least $1-\delta$, for any $z\in[-1,1]$, there exist indices $I\subset\{1,\cdots,n\}$ such that $|z-\sum_{i=1}^{n}M_{i}X_{i}-\sum_{i\in I}(1-M_{i})X_{i}|\leq\varepsilon$ if
+${\color[rgb]{0,0,0}n\geq C\log\left(\frac{C^{\prime}}{\varepsilon\delta}\right),}$ (1)
+where $C,C^{\prime}>0$ are constants, and $C$ depends on $q$.
+Proof.
+Let $M_{1},\cdots,M_{n}\sim\mathrm{Ber}(q)$ and $m:=\sum_{i}M_{i}$. By Hoeffding’s inequality, we have
+${\color[rgb]{0,0,0}\mathbb{P}\left(|m-qn|\leq\varepsilon_{M}n\right)\geq 1-2\exp\left(-\varepsilon_{M}^{2}n\right)}$ (2)
+for $\varepsilon_{M}>0$. Thus, if we set $n\geq\frac{\log(2/\delta)}{\varepsilon_{M}^{2}}$ and $\varepsilon_{M}:=\frac{\min(q,1-q)}{2}$, we have
+$(2q-\beta)n\leq m\leq\beta n$ (3)
+with $\beta:=q+\varepsilon_{M}=\min(\frac{3q}{2},\frac{1+q}{2})\in(0,1)$, with probability at least $1-\delta$.
+In particular, the number of non-vanishing terms in the sum $\sum_{i\in I}(1-M_{i})X_{i}$ is $n-m\geq(1-\beta)n$ as long as each $X_{i}$ is non-zero.
+Now fix $M_{1},\cdots,M_{n}\sim\mathrm{Ber}(q)$ with $(2q-\beta)n\leq m\leq\beta n$.
+The goal is to approximate $z\in[-1,1]$ and $\sum_{i=1}^{n}M_{i}X_{i}$ by subset sum from $\{(1-M_{i})X_{i}:M_{i}=0\}$.
+For simplicity, we split it into two parts:
+$\{(1-M_{i})X_{i}:M_{i}=0\}=\{X^{\prime}_{1},\cdots,X^{\prime}_{n_{1}}\}\cup\{X^{\prime\prime}_{1},\cdots,X^{\prime\prime}_{n_{2}}\},$ (4)
+where the former part is used for approximating $z\in[-1,1]$ and the latter part for approximating $\sum_{i=1}^{n}M_{i}X_{i}$.
+To approximate $z\in[-1,1]$, we can directly apply Corollary 2.5 from Lueker (1998):
+$\mathbb{P}\left(\forall z\in\left[-1,1\right],\exists I\subset\{1,\cdots,n_{1}\}\text{ s.t. }|z-\sum_{i\in I}X^{\prime}_{i}|\leq\varepsilon\right)\geq 1-\delta$ (5)
+whenever $n_{1}\geq 4C\log(8/\varepsilon\delta)$.
+To approximate $\sum_{i=1}^{n}M_{i}X_{i}$, we have to evaluate its norm. By Hoeffding’s inequality on $X_{i}$’s with $M_{i}\not=0$, we have
+$\mathbb{P}\Big{(}\Big{|}\sum_{i=1}^{n}M_{i}X_{i}\Big{|}\leq\alpha m\Big{)}\geq 1-2\exp\left(-\frac{3\alpha^{2}m}{2}\right),$ (6)
+for any fixed $\alpha>0$, whose value will be specified later.
+Thus $\Big{|}\sum_{i=1}^{n}M_{i}X_{i}\Big{|}\leq\alpha m\leq\alpha\beta n$ holds with probability at least $1-\delta$ whenever $m\geq\frac{2\log(2/\delta)}{3\alpha^{2}}$.
+Since $m\leq\beta n$ holds, $n\geq\frac{2\log(2/\delta)}{3\alpha^{2}\beta}$ is enough.
+From the proof of Corollary 3.1 in Lueker (1998), for any $\gamma\in(0,\frac{1}{4})$, we know that
+$\mathbb{P}\left(\forall z\in\left[-\gamma n_{2},\gamma n_{2}\right],\exists I\subset\{1,\cdots,n_{2}\}\text{ s.t. }|z-\sum_{i\in I}X^{\prime\prime}_{i}|\leq\varepsilon\right)$ (7)
+$\geq 1-\delta-2\exp\left(-\frac{(1-4\gamma)^{2}n_{2}-4}{32}\right)$ (8)
+whenever $n_{2}\geq 2C\log(\frac{2}{\varepsilon\delta})$. Thus if
+${\color[rgb]{0,0,0}n_{2}\geq\max\left(2C\log\left(\frac{2}{\varepsilon\delta}\right),\frac{C^{\prime}}{(1-4\gamma)^{2}}\log\left(\frac{C^{\prime\prime}}{\delta}\right)\right)}$ (9)
+holds, we can approximate any $z\in[-\gamma n_{2},\gamma n_{2}]$ by subset sum of $X^{\prime\prime}_{1},\cdots,X^{\prime\prime}_{n_{2}}$ with probability $1-2\delta$.
+Now we assume $(\alpha\beta/\gamma)n\leq n_{2}$.
+For example, if we set $\alpha=\frac{1-\beta}{10-\beta}$ and $\gamma=\frac{1}{5}$, the assumption is satisfied when we split $\{(1-M_{i})X_{i}\}$ by $n_{1}=n_{2}=\frac{(1-\beta)n}{2}$ in (4).
+Then the approximation (7) can be applied to $z=\sum_{i=1}^{n}M_{i}X_{i}$ with high probability since $|\sum_{i=1}^{n}M_{i}X_{i}|\leq\alpha m\leq\alpha\beta n\leq\gamma n_{2}$ holds.
+By combining (5), (6), (9) with the fact that $n=m+n_{1}+n_{2}$, and replacing $\varepsilon$ and $\delta$ with $\varepsilon/2$ and $\delta/6$, respectively, we obtain the desired results. ∎
+By similar arguments as in Lueker (1998) or Pensia et al.
+(2020), we can easily generalize Lemma A.1 to the case where the distribution followed by $X_{i}$ contains the uniform distribution.
+Also, by using the same proof procedure, we can prove the following extension of Lemma A.1:
+Lemma A.2 (Subset-Sum Approximation in Frozen Networks.).
+Let $X_{1},...,X_{n}$ be independent, uniformly distributed random variables so that $X\sim U(-1,1)$, $M_{1},...,M_{n}$ be independent, Bernoulli distributed random variables so that $M_{i}\sim Ber(p)$ for $p\in(0,1)$, and $M^{\prime}_{1},...,M^{\prime}_{n}$ be independent, Bernoulli distributed random variables so that $M^{\prime}_{i}\sim Ber(q)$ for $q\in(0,1)$.
+Let $\varepsilon,\delta\in(0,1)$ be given.
+Then for any $z\in[-1,1]$ there exists a subset $I\subseteq\{1,...,n\}$ so that with probability at least $1-\delta$ we have $\left|z-\sum_{i=1}^{n}M_{i}M^{\prime}_{i}X_{i}-\sum_{i\in I}M_{i}\left(1-M^{\prime}_{i}\right)X_{i}\right|\leq\varepsilon$ if
+$n\geq C\log\left(\frac{C^{\prime}}{\varepsilon\delta}\right),$ (10)
+where $C,C^{\prime}>0$ are constants, and $C$ depends on $p$ and $q$.
+These corollaries seem to be derived by taking $z^{\prime}:=z-\sum_{i=1}^{n}M_{i}X_{i}$ as the target of the Lueker (1998) and Gadhikar et al.
+(2023) corollaries.
+However, it cannot.
+Their corollaries need the target to be deterministic and have constant bounds.
+On the other hand, since $\sum_{i=1}^{n}M_{i}X_{i}$ is stochastic, $z^{\prime}$ is a stochastic value, and the bounds of $z^{\prime}$ are not constant.
+Therefore, $z^{\prime}$ cannot be directly assigned to the conventional corollaries, and our approach is necessary for proofing the subset-sum approximation with frozen variables.
+Finally, by these extended lemmas, we can prove the SLT existence within the frozen network with a sufficiently large width.
+Theorem A.1 (SLT Existence in Frozen Networks).
+Let $\mathcal{D}$ be input data, $f_{T}$ be a target network with depth $L$, and $f_{S}$ be a source network with depth $L+1$, $\bm{\theta}_{S}$ be a parameter of $f_{S}$, and $p_{l},q_{l}\in(0,1)$.
+Assume that these networks have ReLU activation functions, and each element of $\bm{\theta}_{S}$ is uniformly distributed over $[-1,1]$.
+Also assume that $\bm{\theta}_{S}$ is randomly pruned and locked with pruning ratio $p_{l}$ and locking ratio $q_{l}$ for each $l$-th layer, except for 00-th and $1$-st layers.
+Then, with probability at least $1-\delta$, there exists a binary mask $\bm{m}_{S}$ so that each output component $i$ is approximated as $\max_{\bm{x}\in\mathcal{D}}\left\|f_{T,i}(\bm{x})-f_{S,i}(\bm{x};\bm{\theta}_{S}\odot\bm{m}_{S})\right\|<\varepsilon$ if
+$n_{S,l}$ $\geq C_{l}\log\left(\frac{C^{\prime}_{l}}{\min\{\varepsilon_{l},\delta/\rho\}}\right)n_{T,l},$ $(l\geq 1)$ (11)
+$n_{S,0}$ $\geq C_{0}\log\left(\frac{C^{\prime}_{0}}{\min\left\{\varepsilon_{1},\delta/\rho\right\}}\right)d$ $(l=0)$ (12)
+where $C_{l},C_{l}^{\prime}>0$ are constants, and each $C_{l}$ includes $p_{l}$ and $q_{l}$, except for $l=0$.
+Also, $\varepsilon_{l}$ and $\rho$ are as defined in Burkholz (2022a); Gadhikar et al.
+(2023).
+Proof.
+By following the proof procedure of Gadhikar et al.
+(2023) using our Lemma A.2 instead of Lemma 2.4 in Gadhikar et al.
+(2023), we can obtain the desired result as an extension of Theorem 2.5 in Gadhikar et al.
+(2023).
+∎
+Appendix B Additional Experimental Results
+B.1 Accuracy of SLT within Dense ResNet-50
+This section introduces the preliminary experiment used for determining the main experimental setup.
+Figure 7 compares the accuracy of SLTs within a dense ResNet-5050{50}50 source at different SLT sparsity using ImageNet.
+The SLT with $80\%$ sparsity is the most accurate, reaching $66.8\%$ accuracy.
+Figure 7: Accuracy comparison of SLTs of different sparsity within a dense ResNet-5050{50}50 on ImageNet.
+B.2 Accuracy of SLT within Transformer Architectures
+In this section, we search for SLTs within DeiT-S (Touvron et al., 2021), the Transformer architecture for vision tasks, using ImageNet.
+The experimental setup is same as in the DeiT paper, and we use Edge-Popup for finding SLTs.
+As shown in Table 2, compared to the existing SLT search methods, we can find more accurate SLTs from the frozen network, even though the model size is the same or smaller.
+However, the SLT accuracies are significantly lower than that of the weight-trained network.
+While CNN and GNN maintain some SLT accuracy regardless of network pre-processing (i.e., non-frozen, pruning only, or frozen), as shown in Table 1, DeiT shows a significant decrease in SLT accuracy even without freezing.
+It suggests there are some problems specific to the Transformer architecture with respect to SLT.
+As mentioned in Section 5, SLT within Transformer architectures has not yet been established experimentally and theoretically, and investigating this degradation is left for future work.
+Table 2: SLT accuracy comparison in the DeiT-S architecture.
+DeiT-S & ImageNet
+Method
+(Source Net.)
+Weight
+Init.
+Sparsity
+Pruning
+Ratio [%]
+Locking
+Ratio [%]
+Top-1 Test
+Acc. [%]
+Model
+Size [MB]
+Weight Training KU - - - $78.4$ $83.59$
+SLT (Dense) SKC $60$ 00 00 $51.8$ $2.61$
+SLT (Sparse) SKC $60$ $50$ 00 $10.3$ $1.31$
+SLT (Frozen) SKC $60$ $35$ $15$ $40.2$ $1.31$
+SLT (Frozen) SKC $60$ $40$ $20$ $36.9$ $1.04$